### v2

# Introspective Growth: Automatically Advancing LLM Expertise in Technology Judgment

Abstract
Large language models (LLMs) increasingly demonstrate signs of conceptual understanding, yet much of their internal knowledge remains latent, loosely structured, and difficult to access or evaluate. We propose self-questioning as a lightweight and scalable strategy to improve LLMs’ understanding, particularly in domains where success depends on fine-grained semantic distinctions. To evaluate this approach, we introduce a challenging new benchmark of 1.3 million post-2015 computer science patent pairs, characterized by dense technical jargon and strategically complex writing. The benchmark centers on a pairwise differentiation task: can a model distinguish between closely related but substantively different inventions? We show that compared to placebo scientific information, prompting LLMs to generate and answer their own questions - targeting the background knowledge required for the task - significantly improves performance. These self-generated questions and answers activate otherwise underutilized internal knowledge. Allowing LLMs to retrieve answers from external scientific texts further enhances performance, suggesting that model knowledge is compressed and lacks the full richness of the training data. We also find that chain-of-thought prompting and self-questioning converge, though self-questioning remains more effective for improving understanding of technical concepts. Notably, we uncover an asymmetry in prompting: smaller models often generate more fundamental, more open-ended, better-aligned questions for mid-sized models than large models with better understanding do, revealing a new strategy for cross-model collaboration. Altogether, our findings establish self-questioning as both a practical mechanism for automatically improving LLM comprehension, especially in domains with sparse and underrepresented knowledge, and a diagnostic probe of how internal and external knowledge are organized. We open-source our patent dataset at this URL.
1 Introduction
Large language models (LLMs) are increasingly understood not merely as passive text generators, but as systems capable of reasoning about and engaging with complex realities. Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions [25, 46, 51, 61, 60]. Many of the tasks LLMs perform today demand more than simple fact retrieval. Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions [4, 19, 41]. Much of their internal knowledge remains latent, poorly organized, and difficult to access [40, 24]. This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment [23, 24, 26]. In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for a task. This simple strategy of self-questioning, especially when paired with external retrieval, can automatically improve judgment, expose internal knowledge gaps, and activate reasoning processes that not only enhance the model’s own performance but also transfer effectively to larger models, where such reasoning often remains dormant under standard prompting.
Our first contribution is the proposal of a series of behavioral experiments to systematically evaluate self-questioning as a core mechanism for active learning and introspective growth in LLMs. We focus on the model’s understanding of technology concepts, operationalized through a challenging pairwise differentiation task. Instead of assessing understanding through conventional summarization or question-answer (QA) tasks, we present LLMs with pairs of semantically similar but conceptually distinct patent abstracts and ask whether they describe the same invention. A model that truly understands the material must go beyond surface-level similarity and detect subtle functional or mechanistic differences. To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-different computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution. Patents offer a uniquely demanding testbed: they are written in dense legal-technical language, often strategically obfuscated, and their distinction is adversarially verified by expert patent examiners. Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks [36, 64, 12]  [^1] We have more detailed discussions in Appendix A and B.. We embed a structured self-questioning phase between initial exposure and final prediction. Given a pair of closely related patents, the model generates questions targeting both surface-level details and deeper conceptual understanding. It retrieves relevant scientific content to answer its own questions, combining external evidence with internal knowledge. These QA pairs, without verified correctness, are then fed back into the model to inform its final judgment on whether the patents describe the same invention. This setup allows us to assess whether and how self-generated inquiry and targeted retrieval enhance the model’s conceptual understanding (i.e., its accuracy).
Using our framework, we show that prompting LLMs with self-questioning significantly improves their performance on the patent differentiation task, enhancing judgment accuracy, confidence, and response consistency. We then explore three hypotheses to understand what drives this improvement. One possibility is that the questions themselves help the model parse the patent text more effectively by serving as informative cues. Another is that LLMs possess relevant internal knowledge but often fail to access it. A third hypothesis is that even when the model does retrieve relevant knowledge, its internal representations may compress and lose useful information, such that access to external science demonstrably improves performance. To test these possibilities, we design experiments to explore how LLMs’ internal and external knowledge are organized and transferred across model scales, especially when they are sparsely distributed like science, which is our third contribution. We compare the model’s accuracy when given only the questions, when using its questions and self-generated answers, and when using questions and answers derived from science. We find that prompting the model to ask itself questions helps it parse the patent text and improves performance. Feeding LLMs with their self-produced answers also helps, suggesting that LLMs often fail to retrieve relevant internal knowledge. We also find that self-generated answers are less effective than externally sourced ones, however, indicating that hidden compressed capacity remains insufficient for complex judgment tasks, where raw, external scientific knowledge in the form of QA proves beneficial to add richness. We draw a comparison to chain-of-thought (CoT) prompting and find that while they converge, CoT alone fails to recover critical details. This suggests that LLMs lack robust strategies for probing their own internal knowledge. Finally, we investigate how model size influences this process. Interestingly, we find that small models often generate more useful, open-ended, and better-aligned questions for mid-sized models than larger models do, which highlights the importance of matching question complexity to model capacity, as well as a scalable prompting strategy where smaller models help structure the problem in a straightforward way for stronger models to reason through.
2 A Unique Dataset of Hard-to-Distinguish Patents
Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures—it involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies. Drawing from cognitive science and philosophy, a central aspect of understanding is differentiation: recognizing what makes two concepts distinct. This ability is crucial for constructing models of conceptual relationships and is especially vital for causal reasoning, where detecting contrasts and similarities supports counterfactual thinking and explanatory modeling. [^2] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
Because direct questions like “Do you understand X?” often elicit superficial or memorized responses, we need a more robust test of understanding. Instead, we propose a task that evaluates the model’s capacity to discriminate between concepts that are linguistically similar but semantically distinct. Our test centers on pairwise differentiation: the model is presented with two short texts and asked whether they describe the same underlying idea or two distinct ones. These texts are carefully selected to be semantically similar in language and structure — making the task deceptively difficult, and surface-level heuristics ineffective. The model must detect subtle differences in purpose, mechanism, and/or function, and reason about their implications.
We apply this task in the technology domain and focus specifically on patent abstracts. We construct a novel dataset consisting of 1,319,184 pairs of information and computer technology patent abstracts granted by the USPTO after 2015. In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention. We find that the differentiation task requires genuine conceptual linking and functional comparison — capabilities that cannot be achieved simply by decomposing the task into smaller steps (e.g., via CoT reasoning) or matching texts heuristically – making it a strong testbed for evaluating LLM understanding in the technology domain. Several features make technology patents particularly well suited for testing fine-grained semantic distinctions:
(a) Patents are hard to read and understand
(b) Patents are hard to distinguish
Figure 1: Panel (a): Model perplexity of LLaMA-3.1-8B-Instruct on paper and patents abstracts. Panel (b): Histogram of all 1,319,184 patents, showing the distribution of each patent’s maximum cosine similarity to another (distinct) patent.
First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement [53]. These subtle distinctions have been rigorously evaluated by human judges (we will develop this in the third point) and represent precisely the type of contrasts that test whether an LLM can move beyond surface-level similarity and recognize functional difference. We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from [15], pretrained on patent-citation relations. The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
Second, patent documents are often strategically obfuscated, written to disclose the minimum required for legal protection while revealing as little as possible about the underlying invention. As a result, their language tends to be intentionally dense, formal, and legally contorted. Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms [29, 39]. This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper abstracts from ArXiv and 4,000 matched patent abstracts from the same years. Perplexity, while not a direct measure of understanding, reflects the model’s inability to predict the next word in a sequence. Consistently higher perplexity on patent abstracts suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend. An example of paired patent abstracts and a related paper abstract can be seen in Appendix D.
Third, and most importantly, even though patents may seem similar, we know that they are not. Each granted patent in our dataset has undergone an extensive expert evaluation process conducted by patent examiners at the USPTO. This review assesses whether the application satisfies the legal requirements of novelty, non-obviousness, and utility. During this process, examiners conduct detailed prior art searches, compare the proposed invention to existing patents and publications, and determine whether the claimed invention constitutes a meaningful and non-trivial contribution over what already exists. This institutional validation is crucial for our task: it provides a strong, domain-informed guarantee that two different granted patents indeed describe conceptually distinct inventions.
3 Evaluating LLMs’ Understanding of the Technology Domain
We now describe how we operationalize the pairwise differentiation task. We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER [15], a contrastive learning model trained on patent citation relationships. For each patent, we select its most similar—but non-identical—counterpart based on cosine similarity in the embedding space. These pairs are then presented to the LLM, which is asked whether the two abstracts describe the same invention. [^3] See Appendix E for full prompts. For each prompt, the model returns two outputs: a binary label indicating “yes” (1) or “no” (0), and a confidence score on a scale from 0 to 10.
To construct our main measure of understanding, we aggregate predictions using a confidence-weighted voting scheme applied over three independent generations of the same prompt. If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0. This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting [14]. In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases. This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain [14]. Alongside final labels, we also track changes in confidence and measure self-consistency, defined as the proportion of times all three responses yield the same answer. Lower consistency indicates greater uncertainty in model reasoning.
Figure 2: Ask LLMs "Do you think they are the same patent/different patents?", given that they are actually different. All models are Llama 3.1 instruction-tuning versions.
Figure 2 illustrates how models of different sizes perform on the differentiation task within the technology domain. The first panel shows that smaller models, such as LLaMA-8B, struggle to distinguish between semantically similar but distinct patents, classifying over 70% of patent pairs as the same. In contrast, the larger LLaMA-405B model labels only around 10% of these pairs as identical, indicating a stronger ability to detect subtle distinctions. The second and third panels show that the smaller model is also less confident and consistent in its responses (across blue bars). The 8b model’s confidence and self-consistency of wrong answers (blue) is way higher than those of the right answer (orange), again suggesting higher uncertainty and less stable internal representations compared to the larger model. The fourth panel evaluates this phenomenon directly by measuring whether the model gives consistent answers across two question forms. For each patent pair, we ask both “Do you think these patents describe the same invention?” and “Do you think they describe different inventions?” A model with a stable internal representation of the concepts should give logically consistent answers across these two formulations, measured by joint-correct rate, the proportion of pairs where the model gives both correct answers. Deviations from this indicate sensitivity to surface-level wording. The results again show that smaller models are less consistent and more affected by such superficial cues in the phrasing, reinforcing the idea that they lack a robust conceptual understanding of the task.
In the appendix G.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed. All models are more likely to judge the patents as different when asked explicitly whether they are different, even though the content of the question is logically equivalent. We also report additional results for other model families with varying sizes, including Qwen, Mixtral, and GPT-4o-mini. Again, we find that larger models perform better on the differentiation task. Interestingly, we observe consistent response biases in some models: Mixtral models tend to answer “yes” regardless of the question’s phrasing, while Qwen models tend to answer “no.” Based on these findings, we select the LLaMA-3.1-8B Instruct model as the primary focus for our subsequent experiments and use the question “Do you think they are the same patent?” as the default formulation for the pairwise differentiation task.
4 Self-questioning LLMs for Improved Technology Understanding
We begin by examining whether prompting models to ask themselves structured questions can enhance performance and, if so, how. We design two types of self-generated questions. The first, surface-level questions, target conceptual recall (e.g., What is X? What does Y do?), while the second encourages deeper understanding through questions about the integration of concepts, improvement, and the intent underlying invention. We construct the pipeline illustrated in Figure 3. When the model is exposed to a pair of closely related patents, $P_{1}$ and $P_{2}$, it generates six questions $Q_{ij}$ for each, where $i\in{1,2}$ indexes the patent and $j\in{1,\dots,6}$ indexes the question. Three of these are surface-level questions, and three target deeper comprehension. These questions guide the retrieval of relevant scientific knowledge, mimicking how researchers ask clarifying questions to deepen understanding. For external knowledge, we use parsed, pre-processed full texts from arXiv computer science papers. Using Google Scholar, we identify the top 10 papers relevant to the patents, segment each into overlapping 2500-character chunks (with 200-character overlaps), and treat these as candidate knowledge units. We embed both the questions and text chunks using the SPECTER2 model [50] specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question. These are then passed to the LLM as external knowledge $Knowl_{ij}$. The model answers its own questions using both its internal knowledge and the retrieved content, producing $A_{ij}=f_{LLM}(Q_{ij}\mid Knowl_{ij})$. Then the LLM makes a decision by $Decision=Judge(Patent_{1},Patent_{2}\mid\{(Q_{1j},A_{1j})\}_{j=1}^{6},\{(Q_{2j},A_{2j})\}_{j=1}^{6})$ with the previously mentioned confidence-weighted voting schema.
Figure 3: Self-questioning LLMs: The model probes its internals to generate questions, retrieving scientific knowledge or drawing from its own knowledge base to answer them (left). It then uses these structured question-answer pairs to revisit the task under the confidence-weighted voting schema (right).
5 Results
We evaluate whether self-questioning with automatic answer retrieval improves LLMs’ understanding of technology (5.1). We then examine why this strategy works and how it helps us understand the way internal and external knowledge are organized in LLMs – does it aid in activating latent knowledge, offset compressed representations, or primarily rely on external knowledge (5.2). Next, we compare it to CoT prompting to assess whether LLMs can independently construct sufficient background knowledge for high task performance (5.3). We analyze how model size affects question quality and judgment accuracy (5.4).
5.1 Scientific self-questioning works well for boosting LLM understanding
We sample 1000 random pairs and use them as the testbed for demonstration. We gradually add scientific QA pairs from 1-6 for each patent in a pair to the original pair of abstracts, and then have the LLM re-judge. As shown in Figure 4, the baseline represents judge accuracy, confidence, and consistency without any self-questioning or scientific retrieval. The placebo condition uses the same question but adds a randomly selected piece of text from the same paper as the found answers, matched in length to the actual answers. We observe that as more QAs from science are added, judge accuracy significantly increases and becomes better than placebo answers. The confidence of both correct and incorrect answers decreases, but the drop is significantly greater for incorrect answers. Judgments of correct answers also become noticeably more consistent, whereas judgments of incorrect answers become less consistent. Interestingly, even the placebo answer — though merely a random, same-length excerpt from the paper — enhances understanding to some extent. However, unlike actual scientific answers, its effect does not scale with the number of questions meaning there is some initial signal that helps judgment but LLM does not learn more over more rounds of relatively unrelated science. This is likely because: (1) the question alone already facilitates understanding (we will explore this further in the next experiment 5.2), and (2) the placebo text still contains scientific content, which can contribute to comprehension.
Figure 4: More science-QA improves the judge over the placebo answer: For panel (a): $p_{q_{1}}$=0.0868, $p_{q_{2}}$=0.9270, $p_{q_{3}}$=0.0934, $p_{q_{4}}$=0.0000, $p_{q_{5}}$=0.0000, $p_{q_{6}}$=0.0021.
5.2 LLMs do not know what they know
Figure 5: LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA. $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs. scientific answers).
Why does attaching question–answer pairs help? If the model already possesses the relevant knowledge and reliably accesses it when making a judgment, then self-generated questions and answers should have little effect. To explore this, we test whether performance gains stem from organizing internal knowledge, recovering underutilized representations, or introducing new information (Figure 5). We compare three conditions. In the first, the model sees only the questions it generated for each patent, without answers. In the second, it sees both its self-generated questions and answers, without any externally retrieved knowledge (i.e., what it already “knows”, we term this as "self-talk"). In the third, it sees the same questions along with scientific answers retrieved from external sources (the scientific QAs).
Comparing condition one to the raw baseline (blue line), we find that even presenting the model with only its own questions improves judgment, suggesting that structured prompting activates latent knowledge and that the LLM can utilize these textual cues like a “syllabus,” helping it parse complex texts. This effect, however, does not grow with more questions, indicating that the activation is limited. Condition two probes further and examines whether the model already contains the answers in its internal database. The performance improves further by feeding self-talk, and scales with the number of questions, indicating that the model has useful internal knowledge for the task but underutilizes it unless explicitly prompted to retrieve it (orange line). This suggests that the model does not know what it knows in terms of effectively probing internal knowledge  [^4] Another version of “knowing” is more clear-cut – evaluating the validity of an LLMs own claims (true or false) and predicting which questions they will be able to answer correctly [27]. and needs guidance to fully utilize its internal information sources. Finally, condition three (green line) yields the strongest improvement, showing that while the model benefits from questioning and self-generated answers, externally grounded answers still provide a more reliable basis for understanding. The model has often seen the answer content during training, and does not necessarily need to derive it as completely new knowledge through scientific retrieval [56]. In principle, the model already “knows” these facts. (See Appendix F for further experiments of external science acquired through scientific retrieval versus compressed prior known information: We do not find meaningful differences between embedding areas of “content” unleashed through self-talk and retrieved scientific answers. Nevertheless, they are different in other meaningful ways as demonstrated through permutation tests). In summary, model internal representations are highly compressed and do not retain the full richness of information present in the original texts on which they were trained, especially if it is sparsely distributed across the corpus, as with science [52].
5.3 Self-talk and chain-of-thought converge
In our experiments, we explored the “self-talk” QA setup (without scientific retrieval), and we show it works for improving understanding (see Figure 5). This setup can be viewed as a variant of the CoT prompting strategy. To better understand the connection between self-talk and CoT prompting, we conducted two CoT-based experiments: Zero-shot CoT, where we append the instructional prompt “Think step by step, from components to concepts” following the method of [62, 45, 28], and Few-shot CoT, where we provide the model with two annotated examples — one showing a pair of patents that are rephrasing and one showing a pair that is superficially similar but different. We compare performance across four conditions (Figure 6): a single self-talk QA, six rounds of self-talk QA, zero-shot CoT, and few-shot CoT. Our findings reveal that for small-to-mid-sized models, such as LLaMA 8B, zero-shot CoT fails to meaningfully activate reasoning capabilities. While few-shot CoT shows modest improvements, its gains are significantly outpaced by the multi-step self-talk QA approach, as shown in Figure 6 (a). To explore this further we prompted the Llama-8B model to output its few-shot CoT reasoning text across steps explicitly. The model produced $6.70\pm 0.11$ CoT reasoning steps on average. We performed a cumulative-addition experiment where we preserved the original order of the first six CoT steps, and progressively appended QAs one at a time (from 1 to 6) in a random order. At each step $k$, we computed the text similarity between the first $k$ CoT steps and the first $k$ randomized QA texts. As shown in Figure 6 (b), the CoT reasoning and self-talk QAs become semantically aligned and converge across reasoning steps. The initial steps of CoT reasoning are likely more general, but as steps accumulate, CoT becomes increasingly specific and focused, gradually forming stronger semantic connections with the self-talk QA content.
Figure 6: Few-shot CoT and self-talk QAs manifest different performances with self-talk QAs performing better (panel a), but both boost understanding and eventually converge (panel b).
5.4 Questions generated by smaller models are compatible with larger models
As larger models show a higher ability to differentiate, does that also imply they generate better questions? The answer depends on which model is doing the answering. We test two models—LLaMA 3B and 8B—on questions generated by four others: LLaMA 1B, 8B, 70B, and 405B. We select 3B and 8B as answerer models because the 1B model produces low-confidence responses (94.4% with a confidence score of 0), indicating little understanding of content. In contrast, the 70B and 405B models already achieve strong performance without external scientific context, leaving little room for improvement. The 3B and 8B models provide a more informative middle ground: they show some understanding but still benefit from superior prompts.
As shown in Figure 7 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B. This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion. A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales [33]. At the same time, we observe that questions from smaller models often align well with the answering capacity of larger models. This may be because larger models can readily interpret, build on, and give a better answer to smaller models’ questions. Large model questions appear to go beyond the bounds (and dimensionality) of smaller model understanding and answering capacity.
Figure 7: Panels a and b: Questions generated by smaller models are compatible with larger models, but not the converse. Panels c to h: mechanisms. Panels c and d: higher perplexity when answering small models’ questions. Panels e and f: more new information when answering small models’ questions. Panels g and h: smaller models’ questions are more open-ended.
To probe this further, we generate 10,000 question-scientific answer pairs using the 3B and 8B models (the answer model), based on questions from either smaller or larger models relative to the answer models’ own size (smaller or bigger question models), resulting in 40,000 total pairs. We then measure textual similarity between each QA pair and its associated patent abstract. When a larger model answers questions from a smaller model, the resulting QA content is consistently less similar to the original patent abstract, suggesting that small models’ questions bring in more external scientific information to improve the quality, texture, and richness of knowledge beyond the raw patent texts (Figure 7 panels e and f). Small model questions “discover more” external information from the sources they search. We also re-feed the scientific QA pairs into the 3B and 8B models and record their model perplexity. This lower similarity predictably activates more latent reasoning in the models, as reflected in their answer’s higher perplexity (Figure 7 panels c and d). This may stem from how smaller models phrase their questions. With limited capacity, their questions tend to have two important features: (1) Small model questions are more fundamental — 68% of 1B’s questions contain “what is” or “what are” (usage/purpose of components, concepts, and so on), compared to only 44% for 405B. (2) Smaller models tend to ask shorter, less contrastive, and more open-ended questions, potentially yielding higher information entropy (Figure 7 panels g and h). A 1B model’s questions average 16.7 words, compared to 23.0 for a 405B model. We further sampled 200 questions from both smaller and larger question models. The answer model (3B or 8B) was prompted to respond to each question ten times across temperatures from 0.1 to 1, controlling for the answer model’s underlying determinism. We quantify “open-endedness” as the embedding distance of each answer from the mean of all answers (standard deviation). Questions from smaller question models elicit more variable answers.
6 Final Remarks
Limitations: see Appendix H. Conclusions: Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge. While models hold useful internal insights, these are often latent and underused—even CoT prompting cannot fully unlock them. Self-questioning helps activate this hidden compressed capacity, but falls short for complex tasks. Interestingly, the fundamental questions from smaller models often align with the reasoning needs of larger models. This suggests a scalable strategy: use smaller models to structure problems and guide deeper reasoning in stronger models. Impact: our work and the associated new benchmark on patent distinction offer broad insights into how AI can be used to automate the evaluation of scientific and technical contributions.
Acknowledgements
We thank Dr. Ari Holtzman, Sida Li, Jingyang Yi, and Jiazheng Wang from UChicago Data Science Institute and members from UChicago Knowledge Lab for their feedback. This work is completed with the resources provided by the UChicago Research Computing Center. We conducted all experiments using two NVIDIA A100 GPUs (40 GB each) and 150 GB of system RAM.
Appendix
Appendix A Understanding Through Differentiation
Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; [18] it fundamentally involves grasping the relationships that connect concepts and principles within a domain. This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall. Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
Cognitive science, from [16] through [54] and Gentner’s structure-mapping work [35] to recent relational-reasoning research [1], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”. Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’. Without this ability to draw meaningful distinctions, even simple relationships cannot be properly conceptualized. Generalization, conversely, involves identifying commonalities across different instances, abstracting shared properties or underlying structures, and treating distinct items as equivalent for certain purposes, such as belonging to the same category or representing the same concept. These processes are fundamental not just for forming basic concepts, but for constructing the complex relational knowledge that constitutes deep understanding in any domain.
Moreover, the ability to differentiate and generalize remains vital for constructing and refining our understanding of relationships between concepts, especially causal ones, which serve as the backbone of understanding. To explain why something occurs, we must first detect patterns: when does outcome $B$ follow event $A$, and when does it not? This requires the capacity to distinguish $A$ from alternatives like $A^{\prime}$, and to notice that $B$ follows $A$ but not $A^{\prime}$. Such contrasts form the basis for counterfactual reasoning, allowing us to imagine what would happen if everything about $A$ and $A^{\prime}$ were held constant except for one defining feature—if changing that feature flips the outcome, it signals a likely cause. At the same time, we must also recognize when observations are similar in respects that matter. Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes $B$.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims. Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [42, 30, 22]. Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [17].
LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language [57], [58], [38], [8]. Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model [55] that describes and explains how things relate to one another. The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts. Consequently, measuring an LLM’s proficiency in discerning similarity and difference within highly specific concepts and fields offers a powerful proxy for evaluating its understanding of the conceptual foundations and landscape of these fields.
Our measure of understanding draws inspiration from previous benchmarks used to assess language comprehension in earlier generations of language models. GLUE [57], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other. Similarly, our tasks focus on whether models can distinguish subtle conceptual differences. While prior benchmarks often relied on synthetic or crowd-sourced data, e.g [7], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring fine-grained reasoning over technical content.
Appendix B Related Literature
Relations to other benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding. But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [36, 64, 12]. High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension. In particular, current evaluation datasets frequently allow models to exploit statistical associations or training data overlaps to get the right answers without genuine understanding  [^5] https://www.sciencenews.org/article/ai-understanding-reasoning-skill-assess. Even when test content is novel, models can exploit formatting quirks. Alzahrani et al. [2] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions. Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application. Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [3]. Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [9]. In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities. Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [4] or contrast and counterfactual sets [31]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [63]..
Organizing internal and external knowledge: A core challenge in LLM reasoning is how to activate the vast latent knowledge embedded in the model’s billions of parameters. Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [6, 23, 24, 26, 40]. Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted. CoT prompting was one attempt to probe this internal knowledge by encouraging stepwise recall; however, CoT alone cannot guarantee retrieval of all relevant details. The model might skillfully reason with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge. This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [47, 17]. To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources. Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [44]. Lewis et al. [32] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks. Borgeaud et al. [6] introduced the RETRO (Retrieval-Enhanced Transformer) model. A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval. He et al. [20] showed that an LLM with a retrieval step produces more faithful, grounded reasoning – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness. In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries. This is likely because LLMs gain new information during retrieval. LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [49]. By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarity and difference.
Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [21, 13, 10, 11]. Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure. While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.
Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [25, 46, 51, 61, 60]. Instead of only generating direct answers, many contemporary models are encouraged to ask themselves clarifying questions and work through sub-problems, mimicking a Socratic or pedagogical approach. Early prompting techniques like chain-of-thought (CoT) [62] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks. Building on this, researchers introduced methods for the model to generate its own intermediate questions. For example, self-ask prompting [43] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces. Beyond decomposition, recursive self-reflection techniques have emerged to let LLMs iteratively improve their answers. Madaan et al. [34] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds. Shinn et al. [48] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts. Huang et al. [25] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks. Miao et al. present SelfCheck [37], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks. Similarly, techniques like least-to-most prompting [65] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges. Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [59]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work. These introspective or pedagogically-inspired strategies are motivated by human learning: students often learn better by asking themselves questions and reflecting on mistakes. Empirically, such approaches can bolster an LLM’s conceptual grasp and reduce reasoning errors. Nevertheless, there are also limitations and mixed findings. It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [5]. As we demonstrate, large models are much more likely to productively answer small model questions than the converse. Not all tasks benefit uniformly from self-reflection. Li et al. [33] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task. Thus, while self-questioning and introspective prompts represent a powerful paradigm for enhancing LLM reasoning, careful tuning is needed to ensure the model’s “inner dialogue” remains productive and does not overthink or hallucinate.
Appendix C Judging Close Patents VS. Rewritten Patents
(a) Few-shot CoT fails to help distinguish between patents with meaningful technical differences.
(b) Few-shot CoT helps identify similarities across patent abstracts and their rewritten versions.
Figure 8: Few-shot CoT reasoning can distinguish the rephrased abstracts (rewritten by the GPT-o3 model), but cannot distinguish their counterpart (another unique patent).
Following on prior sections that focus on distinguishing deceptively similar patents, in this section, we explore whether and how LLMs are able to merge identical but paraphrased patent ideas. Our approach to measuring LLMs’ ability to generalize and identify the same ideas across distinct articulations is to generate rephrased versions of each patent abstract using the OpenAI ChatGPT o3 model with the prompt “rephrase the patent, but keep its meaning.” We then task the LLaMA-3.1-8B-instruct model with identifying whether a patent abstract and its paraphrase describe the same invention. This was designed to parallel the original distinguishing experiment, in which we randomly sampled 1,000 patents and identified pairs that were most semantically similar. We then asked the LLaMA-3.1-8B-instruct model to determine whether the two abstracts described the same invention, using a human-annotated ground truth indicating that they were different. In raw (zero-shot) judgments, the model performed poorly on both merging and splitting tasks. We then introduced zero-shot CoT prompting (“think step-by-step”) and few-shot CoT, where we supplied examples of both merging (paraphrased) and splitting (semantically similar but distinct) patent pairs. Few-shot CoT substantially improved the model’s accuracy in recognizing paraphrased abstracts as equivalent—but had minimal effect on its ability to distinguish between closely related yet different inventions.
This pattern suggests that CoT reasoning helps the model overcome basic semantic similarity checks and identify paraphrases, likely by encouraging step-by-step alignment at the surface level. Our hypothesis is that the model passes the rephrasing test primarily through pattern matching and semantic overlap, rather than through true conceptual integration. This explains why CoT improves performance on the rephrasing task: the test is relatively easy in the sense that our rewritten version keeps semantic similarity high, such that shallow strategies suffice. In contrast, differentiating between semantically similar but substantively different patents remains hard, even with CoT, because it requires more than aligned phrasing—it demands conceptual depth.
In principle, a more advanced rephrasing model—one that intentionally avoids superficial alignment and instead rewrites concepts in deeper, structure-preserving ways—might push the LLM closer to the more difficult task of distinguishing near matches. Designing such a rephrasing system is far from trivial. Thus, while our current rewrite approach does not fully obscure surface similarity, the model’s success only when the task is easy reinforces the motivation behind our benchmark: LLMs rely heavily on surface-level semantic similarity, rather than demonstrating a deeper understanding of the underlying concepts. In this way, our dataset represents a frontier capability for all contemporary LLMs, even though larger models outperform smaller ones.
Appendix D The Patent Dataset
The patent dataset comprises 1,319,184 patent pairs in total, drawn from five Cooperative Patent Classification (CPC) classes: H05 – Electric techniques not otherwise provided for; H04 – Electric communication technique; G06 – Computing; Calculating; Counting; H03 – Basic electronic circuitry; and G11 – Information storage. Within these, the patents are further categorized into 41 CPC subclasses: [’H05B’, ’H04L’, ’H05H’, ’H05K’, ’G06V’, ’G06N’, ’G06T’, ’H04N’, ’G06F’, ’G06Q’, ’H03M’, ’G06K’, ’H04B’, ’H04R’, ’G11B’, ’G11C’, ’H04M’, ’H03B’, ’H03F’, ’H03G’, ’H03H’, ’H03K’, ’H03L’, ’H04H’, ’H04J’, ’H04K’, ’H04W’, ’H04Q’, ’H04S’, ’H05F’, ’H05G’, ’H03C’, ’H03J’, ’G06E’, ’H03D’, ’G06G’, ’G06M’, ’G06C’, ’H05C’, ’G06J’, ’G06D’]. Out of all matched patent pairs in our dataset, 82.55% share the same CPC class, and 68.10% share the same, more specific CPC subclass. In contrast, only 10.50% of the pairs were granted in the same year.
This is an example patent pair where all LLaMA 3.1 instruct models — 8B, 70B, and 405B - fail to distinguish. We also show a scientific paper abstract to help readers perceive the distinct writing styles between patents and papers.
Patent 1: A wireless device (18) is configured for use in a wireless communication system. The wireless device (18) is configured to receive, from a radio network node (12), a multicast or broadcast transmission (14) and downlink control information indicating a number of repetitions with which the multicast or broadcast transmission (14) is transmitted. The wireless device (18) is also configured to operate according to a negative acknowledgement only feedback scheme in which the wireless device (18) is configured to transmit a negative acknowledgement to the radio network node (12) if decoding of the multicast or broadcast transmission (14) fails using the indicated number of repetitions and to refrain from transmitting a positive acknowledgement to the radio network node (12) if decoding of the multicast or broadcast transmission (14) succeeds.
Patent 2: Methods, systems, and devices for wireless communications are described that support feedback for multicast communications. A user equipment (UE) in a group of UEs may receive multicast information from a base station in a downlink (DL) transmission scheduled by a DL grant. The DL grant may also indicate a set of uplink resources for the group of UEs to use to transmit feedback. The UE may attempt to decode a message containing the multicast information in the DL transmission. If the UE determines the message was successfully decoded, the UE may send no feedback to the base station. If the UE determines the message was not successfully decoded, the UE may transmit a feedback message to the base station to indicate the message was unsuccessfully decoded. The base station may monitor for the feedback message and determine to retransmit the multicast information if the feedback message is received.
An Arxiv paper including the answer from the sampled patent pair: The Third Generation Partnership Project (3GPP) has recently published its Release 16 that includes the first Vehicleto-Everything (V2X) standard based on the 5G New Radio (NR) air interface. 5G NR V2X introduces advanced functionalities on top of the 5G NR air interface to support connected and automated driving use cases with stringent requirements. This paper presents an in-depth tutorial of the 3GPP Release 16 5G NR V2X standard for V2X communications, with a particular focus on the sidelink, since it is the most significant part of 5G NR V2X. The main part of the paper is an in-depth treatment of the key aspects of 5G NR V2X: the physical layer, the resource allocation, the quality of service management, the enhancements introduced to the Uu interface and the mobility management for V2N (Vehicle to Network) communications, as well as the co-existence mechanisms between 5G NR V2X and LTE V2X. We also review the use cases, the system architecture, and describe the evaluation methodology and simulation assumptions for 5G NR V2X. Finally, we provide an outlook on possible 5G NR V2X enhancements, including those identified within Release 17.
Appendix E Prompts
Prompt for Question Generation System Prompt: You are an expert in patent comprehension. Your task is to generate structured questions that assess a reader’s background knowledge necessary to understand a given patent abstract. The questions should focus on foundational concepts, principles, and applications relevant to the patent’s domain without explicitly referencing the patent itself. Ensure the questions are framed generally and test for domain knowledge rather than details specific to the patent. The questions should follow Bloom’s Taxonomy and be categorized into three levels: 1. Remembering: Questions that assess the reader’s ability to recall key technical concepts, definitions, and fundamental principles relevant to the patent’s domain. These should focus on identifying and defining key terms, recognizing fundamental components, and understanding core principles of related technologies. 2. Understanding: Questions that assess the reader’s ability to explain how different elements of similar technologies function and interact. These should encourage summarization, interpretation, and explanation of related technical concepts and their roles in broader systems. 3. Applying: Questions that evaluate the reader’s ability to apply their knowledge by solving problems, making predictions, or considering real-world applications of the broader technological principles underlying the invention. These questions should be scenario-based and encourage practical thinking. Do not directly reference the patent abstract in any question. Instead, ensure all questions test knowledge that would be useful for understanding the patent without explicitly addressing its content. Maintain a structured and logical flow from basic recall to deeper conceptual understanding and practical application. Basic Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Abstract: {patent_summary} Generate a set of {qnum} questions that test a reader’s ability to recall fundamental concepts, key terms, and basic components necessary to understand the given patent abstract. The questions should focus on defining terms, identifying parts, and recognizing the primary function of the described technology. Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }} Conceptual Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Abstract: {patent_summary} Generate a set of {qnum} questions that assess a reader’s comprehension of the given patent abstract by requiring them to explain how different elements of the invention work together. The questions should encourage the reader to summarize the design, describe interactions between components, and interpret the intended improvements of the invention. Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }}
Prompt for Answering Questions with Scientific Knowledge You are a scientific reasoning assistant. Use the following context to answer the question. Focus only on the information provided. Context: // retrieved scientific knowledge {context1} {context2} {context3} Question: {question} 1. Identify and summarize key points from the context relevant to the question. 2. Synthesize those points with your knowledge and try your best to answer the question. Return your answer in the following JSON format: {{ “answer”: “your answer” //your answer to the question }}
Prompt for the Splitting Task Judgment You are a patent judge. Below are two patent abstracts with extra information. This extra information may or may not be useful in assessing the patents. You may choose to use it at your discretion. // This is for placebo answers Your task: judge whether they describe the same patent. Abstract 1: {abstract1} {qa_block1} Abstract 2: {abstract2} {qa_block2} Strictly return your concise answer in the following JSON format: {{ “label”: 1 // 1 if they describe the same patent, 0 if not “confidence”: 10 // your confidence score from 0 to 10, with 10 representing the highest confidence }} // Note: // This applies to the reverse questioning scenario (the major experiments in the paper): // "Do you think they are the same patent?" — when, in fact, the patents are judged to be *different* by human experts. // Invert this setup: Replace "same" with "different", and return 1 if the LLM concludes that the patents are *different*. This yields an alternative version of the experiment (See Section 3 Figure 2 in the paper).
Prompt for Self-talk Answering Try to answer the question. Question: {question} Return your concise answer in the following JSON format: {{ "answer": "your answer" // your answer to the question }}
Few-shot CoT I give you two examples: Example 1 Abstract 1: A method is provided for transmitting control information between a line-switching and a packet-switching communications network. A signaling connection is set up in a packet-switching network in order to transmit signaling packets which form connection-independent control information which relates to at least one service feature in a line-switching communications network, so that at least one service feature of the line-switching communications network may be used in the packet-switching communications network by means of the control information independently of the connection. Abstract 2: A method and apparatus for connecting networks of different types of transmission wherein, in a first network, the data are transmitted in a circuit-switched manner, and in a second network, the data are packet-switched and transmitted connectionlessly on the lower three protocol levels. To be able to perform the signaling in a simple manner, signaling units are used which process signaling messages of the circuit-switched network. Label: The two abstracts are describing different patents according to human expert judgment. Example 2 Abstract 1: A method is provided for transmitting control information between a line-switching and a packet-switching communications network. A signaling connection is set up in a packet-switching network in order to transmit signaling packets which form connection-independent control information which relates to at least one service feature in a line-switching communications network, so that at least one service feature of the line-switching communications network may be used in the packet-switching communications network by means of the control information independently of the connection. Abstract 2: A method is disclosed for transmitting control information between a line-switching and a packet-switching communications network. A signaling connection is established within the packet-switching network to transmit signaling packets that carry connection-independent control information. This control information pertains to at least one service feature of the line-switching communications network, enabling the use of that service feature within the packet-switching communications network independently of any specific connection. Label: The two abstracts describe the same patent, as the second is a rephrasing of the first. // Then attach the Prompt for the Splitting Task Judgment
Appendix F External Knowledge and Compressed Knowledge
In the original paper, we identify a consistent gap between the internal knowledge of LLMs and external scientific knowledge (see Figure 5 in the original paper). We propose two possible explanations for this gap: (1) LLMs function as compressed representations of knowledge. Scientific and technological information, in particular, tends to be sparsely and unevenly distributed in the training corpus, and the compression process significantly degrades the fidelity of such underrepresented knowledge. (2) LLMs acquire new knowledge at inference time — that is, they attempt to reason about or generate responses related to concepts they have never encountered during training but acquired in scientific retrieval.
Both mechanisms may contribute to the observed gap, we argue that the first is the primary cause, for the following reasons:
(a) Panel a
(b) Panel b
Figure 9: Panel a: sample questions. Panel b: scientific retrieval does not introduce more "content-level" new information for LLM to answer these questions.
  • Models tested in the paper — such as LLaMA 3.1, released in July 2024777https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the papers that produced scientific answers had already been posted, and LLaMA 3.2, released around October 2024888https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, when approximately 86% of such papers were available — should have had access to most relevant information.
  • Qualitative inspection of the generated questions suggests that answering them does not require substantial novel information beyond the models’ training. Many questions are straightforward and focus on established concepts or systems. For example, questions involving the TCP Port Service Multiplexer (TCPMUX) relate to long-standing infrastructure, with patents contributing only incremental advancements. As shown in our original paper, 68% of the Llama 1B model’s questions and 44% of the Llama 405B model’s questions include phrases like “what is” or “what are,” indicating the dominance of factual, definitional queries. A sample of questions can be seen in Figure 9 panel a.
  • We further analyzed this by embedding all answers from the scientific sources and the LLM’s (LLaMA 8b) self-talk using the model from [15], specifically pretrained via contrastive learning on patent citation relationships. To visualize and compare the distribution of embedding spaces, we reduce the dimension to 2, estimate the boundary of each region using 5% Gaussian kernel density estimation — that is, we draw contours enclosing the densest 95% of the data points in each embedding set. If LLMs substantially incorporate new scientific content, we would expect their embedding regions to diverge meaningfully from those derived directly from their self-talk. However, as shown in Figure 9 panel b, the embedding regions largely overlap. Interestingly, self-talks display a distinct outlier region, which we suspect may reflect hallucinated content introduced by the LLM.
Despite the similar content - the shared 2D embedding space that represents the greatest variance in high (768)-dimension space, these two types of answers still differ in several meaningful ways. Scientific answers, averaging 62.1$\pm$1.8 words, are possibly richer in detail, drawing more heavily from scientific evidence. In contrast, self-talk answers are notably shorter, averaging 38.2$\pm$0.6 words — only half the length of scientific responses.
Figure 10: The permutation test proves that two types of answers differ significantly in details.
We conducted a permutation test on the centroid distance between the two groups to assess whether self-talk and scientific answers originate from different distributions in embedding space (Figure 10). The test evaluates the null hypothesis that the two groups are exchangeable—that is, their embeddings are drawn from the same underlying distribution. We first computed the Euclidean distance between the centroids of the two groups, which was 1.16 (red line). Then, we randomly permuted the group labels 1,000 times and recalculated the centroid distance for each permutation to generate a null distribution. We find that the observed separation is unlikely under the null hypothesis (reject null, p=0.000). This provides strong statistical evidence that self-talk and scientific answers, although they overlap in coarse-grained content, differ significantly in detail.
Appendix G Results for Majority Voting Schema
G.1 Performances of other LLMs with varying sizes on splitting patents
Figure 11: Ask LLMs "Do you think they are the same patent?", given that they are actually different. Judge using majority voting, and 1 = same. Models are instruction-tuning versions.
Figure 12: Ask LLMs "Do you think they are the same patent?", given that they are actually different. Judge using majority voting, and 1 = same. Models except for GPT-4o-mini are instruction-tuning versions.
Figure 13: Ask LLMs "Do you think they are different patents?", given that they are actually different. Judge using majority voting, and 1 = yes. Models except for GPT-4o-mini are instruction-tuning versions.
G.2 Scientific QAs benefit LLMs’ understanding
(a) QA from science overperforms same questions + placebo answers.
(b) Scientific QA reduces confidence in wrong answers.
(c) Scientific QA reduces wrong answers’ consistency, raises correct answers’ consistency.
Figure 14: Ask Llama 3.1 8B instruct "Do you think they are the same patent?" (splitting task). 1 = yes is the wrong answer.
G.3 LLMs do not know what they have known
Figure 15: Asking Llama 3.1 8B instruct "Do you think they are the same patent?" on actually different patents (1 = wrong answer "yes"). LLM answers using majority voting. It does not know what it has known. These results are statistically significant: Self-talk still helps improve the baseline judge but not as good as scientific self-questioning: $q_{1}$ to $q_{5}$ p=0.000; $q_{6}$ p=0.006.
G.4 Self-talk QAs VS. Chain-of-Thought
Figure 16: Self-talk QAs and CoT prompting manifest different performances.
Figure 17: Self-talk QAs and Few-shot CoT prompting converge.
G.5 Smaller models’ questions help mid-size models more than those from larger models
Figure 18: Llama 3B/8B’s judge with questions proposed by different models with different sizes. Ask “Do you think they are the same patent?”, with the ground truth being *different*. We observe that bigger models’ questions are less useful for the answer models to judge.
Appendix H Discussion of Limitations and Future Work
Several limitations and directions for future work remain. First, we note that certain patent pairs could be effectively indistinguishable based on abstract-level information alone. In such cases, even exhaustive self-questioning by big models fails to resolve ambiguity, suggesting an upper bound on performance determined by input granularity. This highlights the need for richer representations (e.g., claims, diagrams, or specification sections) in future benchmarks. Another, even more intriguing, possibility is that LLMs give the “same” label because they sometimes perceive the underlying similarities between genuinely similar, and potentially follow-on or copying patents. Meanwhile, human judges can also be misled by the storytelling of patent applicants emphasizing nonexistent novelty. In other words, recognizing similarities rather than differences may reflect a more advanced capability for LLMs — a possibility that warrants further investigation.
Second, although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself. Rather than relying on fixed sources or searching questions from scientific databases, LLMs could learn to refine or even generate better retrieval queries via reinforcement learning, active learning, or model-in-the-loop optimization based on science or what they have learned from prior reasoning steps. An agentic framing - where models, like a baby sensing the world, autonomously refine their questions and knowledge, decide whether to retrieve information, evaluate whether to accept or reject what is retrieved, and determine how much knowledge to seek for one question/how many questions to ask - could enable deeper forms of exploration beyond single-pass self-questioning.
Third, the exact nature of “resolution” in distinguishing patents remains underdefined. That is, even when models succeed at effectively splitting, it is not always clear what semantic threshold they are using. Are they resolving based on functionality, application, architecture, or terminology? To what extent are they leveraging scientific knowledge—which aspects, and how relevant is that knowledge to the task? Future work could explicitly model and evaluate these questions to understand what kinds of conceptual distinctions LLMs are truly capable of making by probing their internal, layer-wise conceptual representations and dynamics. It will also be interesting to look into differences in judgment performance across patent classes and categories, as LLM internal knowledge likely varies across fields.
Appendix I Ethics Statement
This research adheres fully to the Code of Ethics. The work does not involve human subjects, crowdsourcing, or sensitive or private data. It does not pose foreseeable risks of misuse, and therefore, no special safeguards are necessary. All external assets, including datasets and pretrained models, are properly credited and used in accordance with their respective licenses. Any new datasets introduced are derived from publicly available sources, anonymized as appropriate, and documented to ensure transparency and reproducibility. Societal impact has been briefly discussed, and the study does not require specific mitigation strategies.

### v3

@@
-Introspective Growth: Automatically Advancing LLM Expertise in Technology Judgment
+Automatically Advancing LLM Expertise in Technology Judgment
 Abstract
-Large language models (LLMs) increasingly demonstrate signs of conceptual understanding, yet much of their internal knowledge remains latent, loosely structured, and difficult to access or evaluate.
-We propose self-questioning as a lightweight and scalable strategy to improve LLMs’ understanding, particularly in domains where success depends on fine-grained semantic distinctions.
+Large language models (LLMs) are rapidly becoming core tools for science, engineering, and innovation.
+Their promise lies not just in remembering facts, but in putting knowledge to work.
+Despite their impressive ability to answer increasingly difficult questions, it remains unclear whether LLMs truly use their knowledge when confronted with new and challenging tasks.
+We address this question with a patent classification task that requires deep conceptual understanding: distinguishing objectively different but semantically similar patents.
 To evaluate this approach, we introduce a challenging new benchmark of 1.3 million post-2015 computer science patent pairs, characterized by dense technical jargon and strategically complex writing.
-The benchmark centers on a pairwise differentiation task: can a model distinguish between closely related but substantively different inventions?
-We show that compared to placebo scientific information, prompting LLMs to generate and answer their own questions - targeting the background knowledge required for the task - significantly improves performance.
-These self-generated questions and answers activate otherwise underutilized internal knowledge.
-Allowing LLMs to retrieve answers from external scientific texts further enhances performance, suggesting that model knowledge is compressed and lacks the full richness of the training data.
-We also find that chain-of-thought prompting and self-questioning converge, though self-questioning remains more effective for improving understanding of technical concepts.
-Notably, we uncover an asymmetry in prompting: smaller models often generate more fundamental, more open-ended, better-aligned questions for mid-sized models than large models with better understanding do, revealing a new strategy for cross-model collaboration.
-Altogether, our findings establish self-questioning as both a practical mechanism for automatically improving LLM comprehension, especially in domains with sparse and underrepresented knowledge, and a diagnostic probe of how internal and external knowledge are organized.
-We open-source our patent dataset at this URL.
+We find that LLMs often fail our benchmark and struggle to distinguish among semantically similar patents.
+To probe this failure, we introduce a novel framework that decomposes model errors into two sources: missing and unused knowledge.
+Our approach asks models to generate clarifying questions to improve their understanding, and then compares three settings: raw performance, self-answered questions, and externally supplied answers.
+This decomposition reveals that LLMs often possess the relevant knowledge internally but fail to deploy it, while a smaller share of errors arises from genuine knowledge gaps.
+We then ask whether the ability of models to construct a task-specific database of questions and answers differs across models.
+We find that smaller models generate simpler, broadly transferable questions, while larger models propose more complex but less generalizable ones.
+This suggests new strategies for combining strengths across models.
+Taken together, our findings highlight a critical limitation of current LLMs and their evaluation: models often know more than they can use.
+By shifting evaluation from recall of static facts to application of dynamic knowledge, our approach provides a more informative lens on model capabilities and opens a path toward building systems that better support technological discovery and innovation.
 1 Introduction
-Large language models (LLMs) are increasingly understood not merely as passive text generators, but as systems capable of reasoning about and engaging with complex realities.
-Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions [25, 46, 51, 61, 60].
-Many of the tasks LLMs perform today demand more than simple fact retrieval.
-Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions [4, 19, 41].
-Much of their internal knowledge remains latent, poorly organized, and difficult to access [40, 24].
-This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment [23, 24, 26].
-In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for a task.
+Large language models (LLMs) are increasingly understood not merely as passive text generators, but systems capable of reasoning about and engaging with complex realities.
+Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions (Huang et al., 2023; Qu et al., 2024; Song et al., 2024; Wang et al., 2023c, b).
+Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023).
+Much of their internal knowledge remains latent, poorly organized, and difficult to access (Pan et al., 2025; Hoang et al., 2023).
+This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment (Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024).
+In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for task completion.
 This simple strategy of self-questioning, especially when paired with external retrieval, can automatically improve judgment, expose internal knowledge gaps, and activate reasoning processes that not only enhance the model’s own performance but also transfer effectively to larger models, where such reasoning often remains dormant under standard prompting.
+A further challenge we highlight is the distinction between lay-in knowledge and working knowledge.
+In humans, once a fact is known, it is typically accessible for reasoning and problem solving: retrieval and use are closely aligned.
+In LLMs, however, the presence of knowledge in parameters does not guarantee its availability in practice.
+Prior work has probed this issue by designing tasks that isolate narrow domains (e.g., math word problems or factual recall) to demonstrate what knowledge LLMs are or are not using (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023).
+By contrast, our framework is situated in a broad technological domain, where the challenge is not the absence of information but the difficulty of activating and applying it in context.
+This generalizable setup, built on a large and diverse patent corpus, goes beyond special-case demonstrations.
+It allows us to investigate systematically how LLMs bridge the gap between latent stored knowledge and the working knowledge required for fine-grained conceptual judgment.
 Our first contribution is the proposal of a series of behavioral experiments to systematically evaluate self-questioning as a core mechanism for active learning and introspective growth in LLMs.
@@
 A model that truly understands the material must go beyond surface-level similarity and detect subtle functional or mechanistic differences.
-To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-different computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution.
-Patents offer a uniquely demanding testbed: they are written in dense legal-technical language, often strategically obfuscated, and their distinction is adversarially verified by expert patent examiners.
-Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks [36, 64, 12]  [^1] We have more detailed discussions in Appendix A and B..
+To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-distinct computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution.
+Patents present a uniquely demanding testbed: they are written in dense, legal-technical language, often strategically obfuscated, and their distinctions are adversarially verified by expert patent examiners.
+Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023)  [^1] We have more detailed discussions in Appendices A and B..
 We embed a structured self-questioning phase between initial exposure and final prediction.
@@
 We also find that self-generated answers are less effective than externally sourced ones, however, indicating that hidden compressed capacity remains insufficient for complex judgment tasks, where raw, external scientific knowledge in the form of QA proves beneficial to add richness.
-We draw a comparison to chain-of-thought (CoT) prompting and find that while they converge, CoT alone fails to recover critical details.
-This suggests that LLMs lack robust strategies for probing their own internal knowledge.
-Finally, we investigate how model size influences this process.
 Interestingly, we find that small models often generate more useful, open-ended, and better-aligned questions for mid-sized models than larger models do, which highlights the importance of matching question complexity to model capacity, as well as a scalable prompting strategy where smaller models help structure the problem in a straightforward way for stronger models to reason through.
 2 A Unique Dataset of Hard-to-Distinguish Patents
-Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures—it involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies.
+Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures.
+It involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies.
 Drawing from cognitive science and philosophy, a central aspect of understanding is differentiation: recognizing what makes two concepts distinct.
@@
 [^2] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
-Because direct questions like “Do you understand X?” often elicit superficial or memorized responses, we need a more robust test of understanding.
+Because direct questions like “what is X?” often elicit superficial or memorized responses, we need a more robust test of understanding.
 Instead, we propose a task that evaluates the model’s capacity to discriminate between concepts that are linguistically similar but semantically distinct.
@@
 In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention.
-We find that the differentiation task requires genuine conceptual linking and functional comparison — capabilities that cannot be achieved simply by decomposing the task into smaller steps (e.g., via CoT reasoning) or matching texts heuristically – making it a strong testbed for evaluating LLM understanding in the technology domain.
-Several features make technology patents particularly well suited for testing fine-grained semantic distinctions:
+We find that the differentiation task requires genuine conceptual linking and functional comparison, making it a strong testbed for evaluating LLM understanding in the technology domain.
+Several features make technology patents particularly well-suited for testing fine-grained semantic distinctions:
 (a) Patents are hard to read and understand
@@
 Panel (b): Histogram of all 1,319,184 patents, showing the distribution of each patent’s maximum cosine similarity to another (distinct) patent.
-First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement [53].
+First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement (Trippe, 2015).
 These subtle distinctions have been rigorously evaluated by human judges (we will develop this in the third point) and represent precisely the type of contrasts that test whether an LLM can move beyond surface-level similarity and recognize functional difference.
-We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from [15], pretrained on patent-citation relations.
+We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from (Ghosh et al., 2024), pretrained on patent-citation relations.
 The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
@@
 As a result, their language tends to be intentionally dense, formal, and legally contorted.
-Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms [29, 39].
+Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms (Kong et al., 2023; Ouellette, 2012).
 This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper abstracts from ArXiv and 4,000 matched patent abstracts from the same years.
@@
 An example of paired patent abstracts and a related paper abstract can be seen in Appendix D.
-Third, and most importantly, even though patents may seem similar, we know that they are not.
+Third, and most importantly, even though patents may seem similar, we know they are not.
 Each granted patent in our dataset has undergone an extensive expert evaluation process conducted by patent examiners at the USPTO.
@@
 We now describe how we operationalize the pairwise differentiation task.
-We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER [15], a contrastive learning model trained on patent citation relationships.
+We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER (Ghosh et al., 2024), a contrastive learning model trained on patent citation relationships.
 For each patent, we select its most similar—but non-identical—counterpart based on cosine similarity in the embedding space.
@@
 If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0.
-This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting [14].
+This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting (Geng et al., 2024).
 In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases.
-This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain [14].
+This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain (Geng et al., 2024).
 Alongside final labels, we also track changes in confidence and measure self-consistency, defined as the proportion of times all three responses yield the same answer.
-Lower consistency indicates greater uncertainty in model reasoning.
+Lower consistency indicates greater uncertainty in model judgment.
 Figure 2: Ask LLMs "Do you think they are the same patent/different patents?", given that they are actually different. All models are Llama 3.1 instruction-tuning versions.
@@
 The results again show that smaller models are less consistent and more affected by such superficial cues in the phrasing, reinforcing the idea that they lack a robust conceptual understanding of the task.
-In the appendix G.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed.
+In the appendix F.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed.
 All models are more likely to judge the patents as different when asked explicitly whether they are different, even though the content of the question is logically equivalent.
@@
 Using Google Scholar, we identify the top 10 papers relevant to the patents, segment each into overlapping 2500-character chunks (with 200-character overlaps), and treat these as candidate knowledge units.
-We embed both the questions and text chunks using the SPECTER2 model [50] specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question.
+We embed both the questions and text chunks using the SPECTER2 model (Singh et al., 2023) specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question.
 These are then passed to the LLM as external knowledge $Knowl_{ij}$.
@@
 5 Results
-We evaluate whether self-questioning with automatic answer retrieval improves LLMs’ understanding of technology (5.1).
-We then examine why this strategy works and how it helps us understand the way internal and external knowledge are organized in LLMs – does it aid in activating latent knowledge, offset compressed representations, or primarily rely on external knowledge (5.2).
-Next, we compare it to CoT prompting to assess whether LLMs can independently construct sufficient background knowledge for high task performance (5.3).
-We analyze how model size affects question quality and judgment accuracy (5.4).
-5.1 Scientific self-questioning works well for boosting LLM understanding
-We sample 1000 random pairs and use them as the testbed for demonstration.
-We gradually add scientific QA pairs from 1-6 for each patent in a pair to the original pair of abstracts, and then have the LLM re-judge.
-As shown in Figure 4, the baseline represents judge accuracy, confidence, and consistency without any self-questioning or scientific retrieval.
-The placebo condition uses the same question but adds a randomly selected piece of text from the same paper as the found answers, matched in length to the actual answers.
-We observe that as more QAs from science are added, judge accuracy significantly increases and becomes better than placebo answers.
-The confidence of both correct and incorrect answers decreases, but the drop is significantly greater for incorrect answers.
-Judgments of correct answers also become noticeably more consistent, whereas judgments of incorrect answers become less consistent.
-Interestingly, even the placebo answer — though merely a random, same-length excerpt from the paper — enhances understanding to some extent.
-However, unlike actual scientific answers, its effect does not scale with the number of questions meaning there is some initial signal that helps judgment but LLM does not learn more over more rounds of relatively unrelated science.
-This is likely because: (1) the question alone already facilitates understanding (we will explore this further in the next experiment 5.2), and (2) the placebo text still contains scientific content, which can contribute to comprehension.
-Figure 4: More science-QA improves the judge over the placebo answer: For panel (a): $p_{q_{1}}$=0.0868, $p_{q_{2}}$=0.9270, $p_{q_{3}}$=0.0934, $p_{q_{4}}$=0.0000, $p_{q_{5}}$=0.0000, $p_{q_{6}}$=0.0021.
+In this section we decompose the error rate into its components - how much of the error is due to missing information and how much is due to the inability to retrieve self-knowledge.
+Section 5.1 shows our main result.
+In the following section, we analyze how model size affects question quality and judgment accuracy (5.2).
 5.1 LLMs Do Not Know What They Know
-Figure 5: LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA.
+(a) LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA.
 $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs.
 scientific answers).
-Why does attaching question–answer pairs help?
-If the model already possesses the relevant knowledge and reliably accesses it when making a judgment, then self-generated questions and answers should have little effect.
-To explore this, we test whether performance gains stem from organizing internal knowledge, recovering underutilized representations, or introducing new information (Figure 5).
-We compare three conditions.
-In the first, the model sees only the questions it generated for each patent, without answers.
-In the second, it sees both its self-generated questions and answers, without any externally retrieved knowledge (i.e., what it already “knows”, we term this as "self-talk").
-In the third, it sees the same questions along with scientific answers retrieved from external sources (the scientific QAs).
-Comparing condition one to the raw baseline (blue line), we find that even presenting the model with only its own questions improves judgment, suggesting that structured prompting activates latent knowledge and that the LLM can utilize these textual cues like a “syllabus,” helping it parse complex texts.
-This effect, however, does not grow with more questions, indicating that the activation is limited.
-Condition two probes further and examines whether the model already contains the answers in its internal database.
-The performance improves further by feeding self-talk, and scales with the number of questions, indicating that the model has useful internal knowledge for the task but underutilizes it unless explicitly prompted to retrieve it (orange line).
-This suggests that the model does not know what it knows in terms of effectively probing internal knowledge  [^4] Another version of “knowing” is more clear-cut – evaluating the validity of an LLMs own claims (true or false) and predicting which questions they will be able to answer correctly [27].
-and needs guidance to fully utilize its internal information sources.
-Finally, condition three (green line) yields the strongest improvement, showing that while the model benefits from questioning and self-generated answers, externally grounded answers still provide a more reliable basis for understanding.
-The model has often seen the answer content during training, and does not necessarily need to derive it as completely new knowledge through scientific retrieval [56].
-In principle, the model already “knows” these facts.
-(See Appendix F for further experiments of external science acquired through scientific retrieval versus compressed prior known information: We do not find meaningful differences between embedding areas of “content” unleashed through self-talk and retrieved scientific answers.
-Nevertheless, they are different in other meaningful ways as demonstrated through permutation tests).
-In summary, model internal representations are highly compressed and do not retain the full richness of information present in the original texts on which they were trained, especially if it is sparsely distributed across the corpus, as with science [52].
-5.3 Self-talk and chain-of-thought converge
-In our experiments, we explored the “self-talk” QA setup (without scientific retrieval), and we show it works for improving understanding (see Figure 5).
-This setup can be viewed as a variant of the CoT prompting strategy.
-To better understand the connection between self-talk and CoT prompting, we conducted two CoT-based experiments: Zero-shot CoT, where we append the instructional prompt “Think step by step, from components to concepts” following the method of [62, 45, 28], and Few-shot CoT, where we provide the model with two annotated examples — one showing a pair of patents that are rephrasing and one showing a pair that is superficially similar but different.
-We compare performance across four conditions (Figure 6): a single self-talk QA, six rounds of self-talk QA, zero-shot CoT, and few-shot CoT.
-Our findings reveal that for small-to-mid-sized models, such as LLaMA 8B, zero-shot CoT fails to meaningfully activate reasoning capabilities.
-While few-shot CoT shows modest improvements, its gains are significantly outpaced by the multi-step self-talk QA approach, as shown in Figure 6 (a).
-To explore this further we prompted the Llama-8B model to output its few-shot CoT reasoning text across steps explicitly.
-The model produced $6.70\pm 0.11$ CoT reasoning steps on average.
-We performed a cumulative-addition experiment where we preserved the original order of the first six CoT steps, and progressively appended QAs one at a time (from 1 to 6) in a random order.
-At each step $k$, we computed the text similarity between the first $k$ CoT steps and the first $k$ randomized QA texts.
-As shown in Figure 6 (b), the CoT reasoning and self-talk QAs become semantically aligned and converge across reasoning steps.
-The initial steps of CoT reasoning are likely more general, but as steps accumulate, CoT becomes increasingly specific and focused, gradually forming stronger semantic connections with the self-talk QA content.
-Figure 6: Few-shot CoT and self-talk QAs manifest different performances with self-talk QAs performing better (panel a), but both boost understanding and eventually converge (panel b).
-5.4 Questions generated by smaller models are compatible with larger models
-As larger models show a higher ability to differentiate, does that also imply they generate better questions?
+(b) Scientific retrieval does not introduce more “content-level” new information for LLM to answer these questions.
+Figure 4: Mechanisms of "LLMs do not know what they know".
+Figure 4(a) presents our initial results from a random sample of 1000 patent pairs, examining how questions and answers affect classification performance.
+The baseline represents the initial accuracy of Llama 8B without any self-questioning or scientific retrieval.
+At this baseline level, Llama 8B incorrectly classifies approximately 70% of patent pairs as identical.
+The blue line demonstrates the impact of adding questions alone.
+This modest improvement suggests that simply posing understanding questions helps the model perform better by providing structure that aids text parsing, though the effect of merely enumerating questions remains limited.
+The orange line reveals that adding self-generated answers significantly improves performance.
+This substantial gain implies that the model already contains useful knowledge for the differentiation task, but this knowledge is not being effectively utilized.
+These results suggest that LLMs do not know what they know—bringing forth relevant information that enhances understanding can dramatically improve performance.
+Finally, the green line shows the improvement achieved by providing retrieved answers.
+This additional performance gain indicates that while the model’s internal knowledge is valuable, it does not match the quality of complete information available in the training data.
+Notably, with six question-answer pairs, performance reaches the same level as the 405B parameter model, effectively eliminating the error rate.
+Importantly, most of the error reduction is achieved through self-generated answers, suggesting that the model contains the most useful information within its internal knowledge representation.
+However, the model lacks the ability to effectively access and apply this knowledge when tackling tasks that require such domain expertise.
+Our results demonstrate that models fail to retrieve the information necessary for the task, and the information they do retrieve is of lower quality than that obtained from additional sources.
+We next explore the factors driving this performance difference, considering two possible explanations.
+First, LLMs function as compressed representations of knowledge but do not preserve complete information.
+Scientific and technological information tends to be sparsely and unevenly distributed in training corpora, and the compression process significantly degrades the fidelity of such underrepresented knowledge.
+Second, LLMs may acquire new knowledge at inference time—meaning the external knowledge we provide is entirely unknown to the LLM and must be learned and incorporated during inference.
+We argue that the first explanation is likely the primary cause of the improvement we observe.
+The models have likely already encountered the data we use: Llama 8B was released in July 2024444https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the retrieved papers that generated scientific answers had already been published.
+However, while the models may have seen these papers during training, this exposure might not have significantly influenced the learned weights.
+To further investigate what type of information is contained in inference-time retrieval versus the model’s internal knowledge, we analyze the answers by embedding all responses from both scientific sources and the LLM’s self-generated answers using the model from (Ghosh et al., 2024).
+To visualize and compare the distribution of embedding spaces, we reduce dimensionality to 2D and estimate boundaries using 5% Gaussian kernel density estimation—drawing contours that enclose the densest 95% of data points in each embedding set.
+If LLMs substantially incorporate new scientific content during inference, we would expect their embedding regions to diverge meaningfully from those derived from self-generated responses.
+However, as shown in Figure 4 panel b, the embedding regions largely overlap.
+Notably, self-generated responses display a distinct outlier region, which we suspect reflects hallucinated content introduced by the LLM.
+Despite occupying similar regions in the 2D embedding space that captures the greatest variance in the high-dimensional (768D) space, these two answer types differ in several meaningful ways.
+Scientific answers average $62.1\pm 1.8$ words and are presumably richer in detail and informativeness, drawing more heavily from scientific evidence.
+In contrast, self-generated answers are notably shorter, averaging only $38.2\pm 0.6$ words—approximately half the length of scientific responses.
+We therefore, interpert the results as suggesting that although the LLMs already have the knowledge needed to answer the question, and obtain better performances - this knowledge is compressed, less detailed and not as rich as the original scientific data.
+5.2 It’s Not Only What to Ask, but Who Asks It
+Self-questioning helps because it aligns the model’s attention with task-relevant concepts.
+This suggests a transferring question: are the most useful questions those that the same model would ask itself, or can questions authored by a different model—smaller or larger—better align the answerer’s knowledge with the task?
+We next test how question–answer sets travel across scale and where misalignment helps or hurts.
 The answer depends on which model is doing the answering.
@@
 We select 3B and 8B as answerer models because the 1B model produces low-confidence responses (94.4% with a confidence score of 0), indicating little understanding of content.
-In contrast, the 70B and 405B models already achieve strong performance without external scientific context, leaving little room for improvement.
+In contrast, the 70B and 405B models already achieve strong performance without an external scientific context, leaving little room for improvement.
 The 3B and 8B models provide a more informative middle ground: they show some understanding but still benefit from superior prompts.
@@
 This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion.
-A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales [33].
+A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales (Li et al., 2024).
 At the same time, we observe that questions from smaller models often align well with the answering capacity of larger models.
@@
 6 Final Remarks
-Limitations: see Appendix H.
-Conclusions: Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge.
-While models hold useful internal insights, these are often latent and underused—even CoT prompting cannot fully unlock them.
+Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge.
+While models hold useful internal insights, these are often latent and underused.
 Self-questioning helps activate this hidden compressed capacity, but falls short for complex tasks.
-Interestingly, the fundamental questions from smaller models often align with the reasoning needs of larger models.
+Interestingly, the fundamental questions from smaller models often align with the needs of larger models.
 This suggests a scalable strategy: use smaller models to structure problems and guide deeper reasoning in stronger models.
 Impact: our work and the associated new benchmark on patent distinction offer broad insights into how AI can be used to automate the evaluation of scientific and technical contributions.
-Acknowledgements
-We thank Dr.
-Ari Holtzman, Sida Li, Jingyang Yi, and Jiazheng Wang from UChicago Data Science Institute and members from UChicago Knowledge Lab for their feedback.
-This work is completed with the resources provided by the UChicago Research Computing Center.
-We conducted all experiments using two NVIDIA A100 GPUs (40 GB each) and 150 GB of system RAM.
 Appendix
 Appendix A Understanding Through Differentiation
-Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; [18] it fundamentally involves grasping the relationships that connect concepts and principles within a domain.
+Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; Grimm [2024] it fundamentally involves grasping the relationships that connect concepts and principles within a domain.
 This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall.
 Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
-Cognitive science, from [16] through [54] and Gentner’s structure-mapping work [35] to recent relational-reasoning research [1], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”.
+Cognitive science, from [Gibson, 1969] through [Tversky, 1977] and Gentner’s structure-mapping work [Markman and Gentner, 2000] to recent relational-reasoning research [Alexander and Dumas, 2018], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”.
 Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’.
@@
 At the same time, we must also recognize when observations are similar in respects that matter.
-Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes $B$.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims.
-Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [42, 30, 22].
-Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [17].
-LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language [57], [58], [38], [8].
-Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model [55] that describes and explains how things relate to one another.
+Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes BB.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims.
+Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [Pearl, 2009, Lake et al., 2017, Heckman and Pinto, 2022].
+Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [Gopnik and Wellman, 2012].
+LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language Wang et al.
+[2018], Wang et al.
+[2019], OpenAI [2023], Brown et al.
+[2020].
+Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model Vafa et al.
+[2024] that describes and explains how things relate to one another.
 The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts.
@@
 Our measure of understanding draws inspiration from previous benchmarks used to assess language comprehension in earlier generations of language models.
-GLUE [57], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other.
+GLUE Wang et al.
+[2018], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other.
 Similarly, our tasks focus on whether models can distinguish subtle conceptual differences.
-While prior benchmarks often relied on synthetic or crowd-sourced data, e.g [7], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring fine-grained reasoning over technical content.
+While prior benchmarks often relied on synthetic or crowd-sourced data, e.g.
+Bowman et al.
+[2015], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring a fine-grained understanding of technical content.
 Appendix B Related Literature
 Relations to other benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding.
-But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [36, 64, 12].
+But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [McIntosh et al., 2024, Wu et al., 2024, Davis, 2023].
 High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension.
@@
 Alzahrani et al.
-[2] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions.
+[Alzahrani et al., 2024] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions.
 Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application.
-Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [3].
-Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [9].
+Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [Arakelyan et al., 2024].
+Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [Burnell et al., 2023].
 In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities.
-Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [4] or contrast and counterfactual sets [31]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [63]..
-Organizing internal and external knowledge: A core challenge in LLM reasoning is how to activate the vast latent knowledge embedded in the model’s billions of parameters.
-Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [6, 23, 24, 26, 40].
+Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [Asthana et al., 2024] or contrast and counterfactual sets [Lewis and Mitchell, 2024]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [West et al., 2024]..
+Organizing internal and external knowledge: A core challenge in LLM understanding is how to activate the vast latent knowledge embedded in the model’s billions of parameters.
+Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [Borgeaud et al., 2022, Hoang et al., , Hoang et al., 2023, Jaiswal et al., 2024, Pan et al., 2025].
 Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted.
-CoT prompting was one attempt to probe this internal knowledge by encouraging stepwise recall; however, CoT alone cannot guarantee retrieval of all relevant details.
-The model might skillfully reason with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge.
-This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [47, 17].
+The model might skillfully "reason" with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge.
+This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [Schulhoff et al., 2024, Gopnik and Wellman, 2012].
 To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources.
-Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [44].
+Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [Prince et al., 2024].
 Lewis et al.
-[32] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks.
+[Lewis et al., 2020] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks.
 Borgeaud et al.
-[6] introduced the RETRO (Retrieval-Enhanced Transformer) model.
+[Borgeaud et al., 2022] introduced the RETRO (Retrieval-Enhanced Transformer) model.
 A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval.
 He et al.
-[20] showed that an LLM with a retrieval step produces more faithful, grounded reasoning – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness.
+[He et al., 2022] showed that an LLM with a retrieval step produces more faithful, grounded understanding – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness.
 In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries.
 This is likely because LLMs gain new information during retrieval.
-LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [49].
-By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarity and difference.
-Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [21, 13, 10, 11].
+LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [Shumailov et al., 2024].
+By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarities and differences.
+Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [He et al., 2020, Gao et al., 2021, Chen et al., 2020, Cohan et al., 2020].
 Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure.
 While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.
-Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [25, 46, 51, 61, 60].
+Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [Huang et al., 2023, Qu et al., 2024, Song et al., 2024, Wang et al., 2023c, b].
 Instead of only generating direct answers, many contemporary models are encouraged to ask themselves clarifying questions and work through sub-problems, mimicking a Socratic or pedagogical approach.
-Early prompting techniques like chain-of-thought (CoT) [62] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks.
+Early prompting techniques like chain-of-thought (CoT) [Wei et al., 2022] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks.
 Building on this, researchers introduced methods for the model to generate its own intermediate questions.
-For example, self-ask prompting [43] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces.
+For example, self-ask prompting [Press et al., 2022] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces.
 Beyond decomposition, recursive self-reflection techniques have emerged to let LLMs iteratively improve their answers.
 Madaan et al.
-[34] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds.
+[Madaan et al., 2023] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds.
 Shinn et al.
-[48] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts.
+[Shinn et al., 2023] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts.
 Huang et al.
-[25] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks.
+[Huang et al., 2023] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks.
 Miao et al.
-present SelfCheck [37], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks.
-Similarly, techniques like least-to-most prompting [65] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges.
-Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [59]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work.
+present SelfCheck [Miao et al., 2024], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks.
+Similarly, techniques like least-to-most prompting [Zhou et al., 2022] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges.
+Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [Wang et al., 2023a]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work.
 These introspective or pedagogically-inspired strategies are motivated by human learning: students often learn better by asking themselves questions and reflecting on mistakes.
@@
 Nevertheless, there are also limitations and mixed findings.
-It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [5].
+It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [Balepur et al., 2025].
 As we demonstrate, large models are much more likely to productively answer small model questions than the converse.
@@
 Li et al.
-[33] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task.
+[Li et al., 2024] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task.
 Thus, while self-questioning and introspective prompts represent a powerful paradigm for enhancing LLM reasoning, careful tuning is needed to ensure the model’s “inner dialogue” remains productive and does not overthink or hallucinate.
@@
 Prompt for Self-talk Answering Try to answer the question. Question: {question} Return your concise answer in the following JSON format: {{ "answer": "your answer" // your answer to the question }}
-Few-shot CoT I give you two examples: Example 1 Abstract 1: A method is provided for transmitting control information between a line-switching and a packet-switching communications network.
-A signaling connection is set up in a packet-switching network in order to transmit signaling packets which form connection-independent control information which relates to at least one service feature in a line-switching communications network, so that at least one service feature of the line-switching communications network may be used in the packet-switching communications network by means of the control information independently of the connection.
-Abstract 2: A method and apparatus for connecting networks of different types of transmission wherein, in a first network, the data are transmitted in a circuit-switched manner, and in a second network, the data are packet-switched and transmitted connectionlessly on the lower three protocol levels.
-To be able to perform the signaling in a simple manner, signaling units are used which process signaling messages of the circuit-switched network.
-Label: The two abstracts are describing different patents according to human expert judgment.
-Example 2 Abstract 1: A method is provided for transmitting control information between a line-switching and a packet-switching communications network.
-A signaling connection is set up in a packet-switching network in order to transmit signaling packets which form connection-independent control information which relates to at least one service feature in a line-switching communications network, so that at least one service feature of the line-switching communications network may be used in the packet-switching communications network by means of the control information independently of the connection.
-Abstract 2: A method is disclosed for transmitting control information between a line-switching and a packet-switching communications network.
-A signaling connection is established within the packet-switching network to transmit signaling packets that carry connection-independent control information.
-This control information pertains to at least one service feature of the line-switching communications network, enabling the use of that service feature within the packet-switching communications network independently of any specific connection.
-Label: The two abstracts describe the same patent, as the second is a rephrasing of the first.
-// Then attach the Prompt for the Splitting Task Judgment
-Appendix F External Knowledge and Compressed Knowledge
-In the original paper, we identify a consistent gap between the internal knowledge of LLMs and external scientific knowledge (see Figure 5 in the original paper).
-We propose two possible explanations for this gap: (1) LLMs function as compressed representations of knowledge.
-Scientific and technological information, in particular, tends to be sparsely and unevenly distributed in the training corpus, and the compression process significantly degrades the fidelity of such underrepresented knowledge.
-(2) LLMs acquire new knowledge at inference time — that is, they attempt to reason about or generate responses related to concepts they have never encountered during training but acquired in scientific retrieval.
-Both mechanisms may contribute to the observed gap, we argue that the first is the primary cause, for the following reasons:
-(a) Panel a
-(b) Panel b
-Figure 9: Panel a: sample questions. Panel b: scientific retrieval does not introduce more "content-level" new information for LLM to answer these questions.
-  • Models tested in the paper — such as LLaMA 3.1, released in July 2024777https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the papers that produced scientific answers had already been posted, and LLaMA 3.2, released around October 2024888https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, when approximately 86% of such papers were available — should have had access to most relevant information.
-  • Qualitative inspection of the generated questions suggests that answering them does not require substantial novel information beyond the models’ training.
-Many questions are straightforward and focus on established concepts or systems.
-For example, questions involving the TCP Port Service Multiplexer (TCPMUX) relate to long-standing infrastructure, with patents contributing only incremental advancements.
-As shown in our original paper, 68% of the Llama 1B model’s questions and 44% of the Llama 405B model’s questions include phrases like “what is” or “what are,” indicating the dominance of factual, definitional queries.
-A sample of questions can be seen in Figure 9 panel a.
-  • We further analyzed this by embedding all answers from the scientific sources and the LLM’s (LLaMA 8b) self-talk using the model from [15], specifically pretrained via contrastive learning on patent citation relationships.
-To visualize and compare the distribution of embedding spaces, we reduce the dimension to 2, estimate the boundary of each region using 5% Gaussian kernel density estimation — that is, we draw contours enclosing the densest 95% of the data points in each embedding set.
-If LLMs substantially incorporate new scientific content, we would expect their embedding regions to diverge meaningfully from those derived directly from their self-talk.
-However, as shown in Figure 9 panel b, the embedding regions largely overlap.
-Interestingly, self-talks display a distinct outlier region, which we suspect may reflect hallucinated content introduced by the LLM.
-Despite the similar content - the shared 2D embedding space that represents the greatest variance in high (768)-dimension space, these two types of answers still differ in several meaningful ways.
-Scientific answers, averaging 62.1$\pm$1.8 words, are possibly richer in detail, drawing more heavily from scientific evidence.
-In contrast, self-talk answers are notably shorter, averaging 38.2$\pm$0.6 words — only half the length of scientific responses.
-Figure 10: The permutation test proves that two types of answers differ significantly in details.
-We conducted a permutation test on the centroid distance between the two groups to assess whether self-talk and scientific answers originate from different distributions in embedding space (Figure 10).
-The test evaluates the null hypothesis that the two groups are exchangeable—that is, their embeddings are drawn from the same underlying distribution.
-We first computed the Euclidean distance between the centroids of the two groups, which was 1.16 (red line).
-Then, we randomly permuted the group labels 1,000 times and recalculated the centroid distance for each permutation to generate a null distribution.
-We find that the observed separation is unlikely under the null hypothesis (reject null, p=0.000).
-This provides strong statistical evidence that self-talk and scientific answers, although they overlap in coarse-grained content, differ significantly in detail.
-Appendix G Results for Majority Voting Schema
-G.1 Performances of other LLMs with varying sizes on splitting patents
+Appendix F Results for Majority Voting Schema
+F.1 Performances of other LLMs with varying sizes on splitting patents
 Figure 7: Ask LLMs "Do you think they are the same patent?", given that they are actually different. Judge using majority voting, and 1 = same. Models are instruction-tuning versions.
@@
 Models except for GPT-4o-mini are instruction-tuning versions.
-G.2 Scientific QAs benefit LLMs’ understanding
-(a) QA from science overperforms same questions + placebo answers.
-(b) Scientific QA reduces confidence in wrong answers.
-(c) Scientific QA reduces wrong answers’ consistency, raises correct answers’ consistency.
-Figure 14: Ask Llama 3.1 8B instruct "Do you think they are the same patent?" (splitting task). 1 = yes is the wrong answer.
-G.3 LLMs do not know what they have known
+F.2 LLMs do not know what they have known
 Figure 10: Asking Llama 3.1 8B instruct "Do you think they are the same patent?" on actually different patents (1 = wrong answer "yes").
@@
 These results are statistically significant: Self-talk still helps improve the baseline judge but not as good as scientific self-questioning: $q_{1}$ to $q_{5}$ p=0.000; $q_{6}$ p=0.006.
-G.4 Self-talk QAs VS. Chain-of-Thought
-Figure 16: Self-talk QAs and CoT prompting manifest different performances.
-Figure 17: Self-talk QAs and Few-shot CoT prompting converge.
-G.5 Smaller models’ questions help mid-size models more than those from larger models
+F.3 Smaller models’ questions help mid-size models more than those from larger models
 Figure 11: Llama 3B/8B’s judge with questions proposed by different models with different sizes.
@@
 We observe that bigger models’ questions are less useful for the answer models to judge.
-Appendix H Discussion of Limitations and Future Work
+Appendix G Discussion of Limitations and Future Work
 Several limitations and directions for future work remain.
-First, we note that certain patent pairs could be effectively indistinguishable based on abstract-level information alone.
-In such cases, even exhaustive self-questioning by big models fails to resolve ambiguity, suggesting an upper bound on performance determined by input granularity.
-This highlights the need for richer representations (e.g., claims, diagrams, or specification sections) in future benchmarks.
-Another, even more intriguing, possibility is that LLMs give the “same” label because they sometimes perceive the underlying similarities between genuinely similar, and potentially follow-on or copying patents.
-Meanwhile, human judges can also be misled by the storytelling of patent applicants emphasizing nonexistent novelty.
-In other words, recognizing similarities rather than differences may reflect a more advanced capability for LLMs — a possibility that warrants further investigation.
-Second, although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself.
+Although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself.
 Rather than relying on fixed sources or searching questions from scientific databases, LLMs could learn to refine or even generate better retrieval queries via reinforcement learning, active learning, or model-in-the-loop optimization based on science or what they have learned from prior reasoning steps.
 An agentic framing - where models, like a baby sensing the world, autonomously refine their questions and knowledge, decide whether to retrieve information, evaluate whether to accept or reject what is retrieved, and determine how much knowledge to seek for one question/how many questions to ask - could enable deeper forms of exploration beyond single-pass self-questioning.
-Third, the exact nature of “resolution” in distinguishing patents remains underdefined.
+The exact nature of “resolution” in distinguishing patents remains underdefined.
 That is, even when models succeed at effectively splitting, it is not always clear what semantic threshold they are using.
 Are they resolving based on functionality, application, architecture, or terminology?
-To what extent are they leveraging scientific knowledge—which aspects, and how relevant is that knowledge to the task?
 Future work could explicitly model and evaluate these questions to understand what kinds of conceptual distinctions LLMs are truly capable of making by probing their internal, layer-wise conceptual representations and dynamics.
 It will also be interesting to look into differences in judgment performance across patent classes and categories, as LLM internal knowledge likely varies across fields.
-Appendix I Ethics Statement
-This research adheres fully to the Code of Ethics.
-The work does not involve human subjects, crowdsourcing, or sensitive or private data.
-It does not pose foreseeable risks of misuse, and therefore, no special safeguards are necessary.
-All external assets, including datasets and pretrained models, are properly credited and used in accordance with their respective licenses.
-Any new datasets introduced are derived from publicly available sources, anonymized as appropriate, and documented to ensure transparency and reproducibility.
-Societal impact has been briefly discussed, and the study does not require specific mitigation strategies.

### v4

@@
-Automatically Advancing LLM Expertise in Technology Judgment
+The Missing vs. Unused Knowledge Hypothesis for Language Model Bottlenecks in Patent Understanding
 Abstract
-Large language models (LLMs) are rapidly becoming core tools for science, engineering, and innovation.
-Their promise lies not just in remembering facts, but in putting knowledge to work.
-Despite their impressive ability to answer increasingly difficult questions, it remains unclear whether LLMs truly use their knowledge when confronted with new and challenging tasks.
-We address this question with a patent classification task that requires deep conceptual understanding: distinguishing objectively different but semantically similar patents.
-To evaluate this approach, we introduce a challenging new benchmark of 1.3 million post-2015 computer science patent pairs, characterized by dense technical jargon and strategically complex writing.
-We find that LLMs often fail our benchmark and struggle to distinguish among semantically similar patents.
-To probe this failure, we introduce a novel framework that decomposes model errors into two sources: missing and unused knowledge.
-Our approach asks models to generate clarifying questions to improve their understanding, and then compares three settings: raw performance, self-answered questions, and externally supplied answers.
-This decomposition reveals that LLMs often possess the relevant knowledge internally but fail to deploy it, while a smaller share of errors arises from genuine knowledge gaps.
-We then ask whether the ability of models to construct a task-specific database of questions and answers differs across models.
-We find that smaller models generate simpler, broadly transferable questions, while larger models propose more complex but less generalizable ones.
-This suggests new strategies for combining strengths across models.
-Taken together, our findings highlight a critical limitation of current LLMs and their evaluation: models often know more than they can use.
-By shifting evaluation from recall of static facts to application of dynamic knowledge, our approach provides a more informative lens on model capabilities and opens a path toward building systems that better support technological discovery and innovation.
+While large language models (LLMs) excel at factual recall, the real challenge lies in knowledge application.
+A gap persists between their ability to answer complex questions and their effectiveness in performing tasks that require that knowledge.
+We investigate this gap using a patent classification problem that requires deep conceptual understanding to distinguish semantically similar but objectively different patents written in dense, strategic technical language.
+We find that LLMs often struggle with this distinction.
+To diagnose the source of these failures, we introduce a framework that decomposes model errors into two categories: missing knowledge and unused knowledge.
+Our method prompts models to generate clarifying questions and compares three settings – raw performance, self-answered questions that activate internal knowledge, and externally provided answers that supply missing knowledge (if any).
+We show that most errors stem from failures to deploy existing knowledge rather than from true knowledge gaps.
+We also examine how models differ in constructing task-specific question–answer databases.
+Smaller models tend to generate simpler questions that they, and other models, can retrieve and use effectively, whereas larger models produce more complex questions that are less effective, suggesting complementary strengths across model scales.
+Together, our findings highlight that shifting evaluation from static fact recall to dynamic knowledge application offers a more informative view of model capabilities.
+The Missing vs. Unused Knowledge Hypothesis for Language Model Bottlenecks in Patent Understanding
 1 Introduction
-Large language models (LLMs) are increasingly understood not merely as passive text generators, but systems capable of reasoning about and engaging with complex realities.
-Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions (Huang et al., 2023; Qu et al., 2024; Song et al., 2024; Wang et al., 2023c, b).
-Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023).
-Much of their internal knowledge remains latent, poorly organized, and difficult to access (Pan et al., 2025; Hoang et al., 2023).
-This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment (Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024).
-In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for task completion.
-This simple strategy of self-questioning, especially when paired with external retrieval, can automatically improve judgment, expose internal knowledge gaps, and activate reasoning processes that not only enhance the model’s own performance but also transfer effectively to larger models, where such reasoning often remains dormant under standard prompting.
-A further challenge we highlight is the distinction between lay-in knowledge and working knowledge.
-In humans, once a fact is known, it is typically accessible for reasoning and problem solving: retrieval and use are closely aligned.
-In LLMs, however, the presence of knowledge in parameters does not guarantee its availability in practice.
-Prior work has probed this issue by designing tasks that isolate narrow domains (e.g., math word problems or factual recall) to demonstrate what knowledge LLMs are or are not using (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023).
-By contrast, our framework is situated in a broad technological domain, where the challenge is not the absence of information but the difficulty of activating and applying it in context.
-This generalizable setup, built on a large and diverse patent corpus, goes beyond special-case demonstrations.
-It allows us to investigate systematically how LLMs bridge the gap between latent stored knowledge and the working knowledge required for fine-grained conceptual judgment.
-Our first contribution is the proposal of a series of behavioral experiments to systematically evaluate self-questioning as a core mechanism for active learning and introspective growth in LLMs.
-We focus on the model’s understanding of technology concepts, operationalized through a challenging pairwise differentiation task.
-Instead of assessing understanding through conventional summarization or question-answer (QA) tasks, we present LLMs with pairs of semantically similar but conceptually distinct patent abstracts and ask whether they describe the same invention.
-A model that truly understands the material must go beyond surface-level similarity and detect subtle functional or mechanistic differences.
-To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-distinct computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution.
-Patents present a uniquely demanding testbed: they are written in dense, legal-technical language, often strategically obfuscated, and their distinctions are adversarially verified by expert patent examiners.
-Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023)  [^1] We have more detailed discussions in Appendices A and B..
-We embed a structured self-questioning phase between initial exposure and final prediction.
-Given a pair of closely related patents, the model generates questions targeting both surface-level details and deeper conceptual understanding.
-It retrieves relevant scientific content to answer its own questions, combining external evidence with internal knowledge.
-These QA pairs, without verified correctness, are then fed back into the model to inform its final judgment on whether the patents describe the same invention.
-This setup allows us to assess whether and how self-generated inquiry and targeted retrieval enhance the model’s conceptual understanding (i.e., its accuracy).
-Using our framework, we show that prompting LLMs with self-questioning significantly improves their performance on the patent differentiation task, enhancing judgment accuracy, confidence, and response consistency.
-We then explore three hypotheses to understand what drives this improvement.
-One possibility is that the questions themselves help the model parse the patent text more effectively by serving as informative cues.
-Another is that LLMs possess relevant internal knowledge but often fail to access it.
-A third hypothesis is that even when the model does retrieve relevant knowledge, its internal representations may compress and lose useful information, such that access to external science demonstrably improves performance.
-To test these possibilities, we design experiments to explore how LLMs’ internal and external knowledge are organized and transferred across model scales, especially when they are sparsely distributed like science, which is our third contribution.
-We compare the model’s accuracy when given only the questions, when using its questions and self-generated answers, and when using questions and answers derived from science.
-We find that prompting the model to ask itself questions helps it parse the patent text and improves performance.
-Feeding LLMs with their self-produced answers also helps, suggesting that LLMs often fail to retrieve relevant internal knowledge.
-We also find that self-generated answers are less effective than externally sourced ones, however, indicating that hidden compressed capacity remains insufficient for complex judgment tasks, where raw, external scientific knowledge in the form of QA proves beneficial to add richness.
-Interestingly, we find that small models often generate more useful, open-ended, and better-aligned questions for mid-sized models than larger models do, which highlights the importance of matching question complexity to model capacity, as well as a scalable prompting strategy where smaller models help structure the problem in a straightforward way for stronger models to reason through.
-2 A Unique Dataset of Hard-to-Distinguish Patents
-Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures.
-It involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies.
-Drawing from cognitive science and philosophy, a central aspect of understanding is differentiation: recognizing what makes two concepts distinct.
-This ability is crucial for constructing models of conceptual relationships and is especially vital for causal reasoning, where detecting contrasts and similarities supports counterfactual thinking and explanatory modeling.
+Large Language Models (LLMs) are increasingly viewed not merely as passive text generators, but as systems that can reason about and engage with complex realities.
+Yet despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023).
+We hypothesize that this difficulty reflects not a lack of knowledge, but limits in how knowledge is represented and accessed: much of what LLMs encode remains latent, poorly organized, and difficult to retrieve in knowledge-intensive settings (Pan et al., 2025; Hoang et al., 2023).
+This creates a fundamental tension: although LLMs can appear to “know everything,” their knowledge is stored in a highly compressed form that can hinder activation, recall, and information-quality assessment (Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024).
+More broadly, the gap between the knowledge an LLM contains and what it reliably deploys at inference time is central to evaluating its competence.
+In humans, knowledge that is learned is typically accessible for reasoning and problem-solving; in LLMs, by contrast, having information in parameters does not guarantee its availability when needed.
+Prior work has examined this gap using narrowly scoped probes, such as math word problems or factual recall, to test what models do or do not use (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023), but such tasks often miss the holistic, integrated nature of conceptual understanding.
+We refer to these two forms of knowledge as lay in knowledge, knowledge the model can produce when directly prompted, and working knowledge, the knowledge it actually brings to bear in the course of solving a task.
+To quantify the contribution of each type of knowledge to task performance, we design a conceptual understanding test that requires deploying knowledge dynamically rather than merely stating it.
+Understanding is a core cognitive capacity underlying most knowledge-intensive behavior: it goes beyond recalling facts or executing procedures, and instead involves grasping structural links and hierarchies that support transfer and generalization.
+Without some levels of understanding, both humans and LLMs may succeed on familiar patterns yet fail under distributional shift, novel recombinations of known components, or adversarially phrased problems.
+We therefore treat understanding not as an optional “nice-to-have,” but as a prerequisite for turning stored information into usable working knowledge.
+Our testbed is patent understanding, where success depends on fine-grained access to mechanisms, purposes, and constraints.
 [^1] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
-Because direct questions like “what is X?” often elicit superficial or memorized responses, we need a more robust test of understanding.
-Instead, we propose a task that evaluates the model’s capacity to discriminate between concepts that are linguistically similar but semantically distinct.
-Our test centers on pairwise differentiation: the model is presented with two short texts and asked whether they describe the same underlying idea or two distinct ones.
-These texts are carefully selected to be semantically similar in language and structure — making the task deceptively difficult, and surface-level heuristics ineffective.
-The model must detect subtle differences in purpose, mechanism, and/or function, and reason about their implications.
-We apply this task in the technology domain and focus specifically on patent abstracts.
-We construct a novel dataset consisting of 1,319,184 pairs of information and computer technology patent abstracts granted by the USPTO after 2015.
+Drawing from cognitive science and philosophy, we posit that a key, observable facet of understanding is differentiation – the ability to recognize what makes two concepts distinct.
+[^2] We provide a more detailed discussion in Appendices A and B.
+Accordingly, we evaluate in this paper models’ capacity to discriminate between complex technological concepts that are linguistically similar yet functionally distinct.
+In our task, a model is presented with two texts (patent descriptions) and asked whether they describe the same underlying idea or two distinct ones.
+These texts are carefully selected to be semantically similar in language and structure, making the task deceptively difficult and surface-level heuristics ineffective.
+To succeed, the model must detect subtle differences in aspects such as purpose and mechanism.
+We choose patents as our primary testbed because they present a uniquely demanding environment; they are written in dense, legal-technical language, often strategically obfuscated, and their distinctions are adversarially verified by human patent examiners.
+This task requires fine-grained conceptual discrimination that goes well beyond the pattern matching (in strategically confusing texts), imitation, or memorization that often bypasses existing benchmarks (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023).
+Our results show that while frontier models display some competence, weaker models consistently miss key technological nuances.
+Building on this, we quantify whether failures on knowledge-intensive understanding tasks stem from missing knowledge or from having “lay-in” knowledge that models cannot reliably use.
+To measure the gap between working knowledge and lay-in knowledge, we prompt models to generate clarification questions to improve understanding about patents and compare baseline performance to two conditions: (i) questions and answers generated by the model itself, and (ii) questions and answers generated using external scientific information.
+We find that understanding failures are driven primarily by poor retrieval of lay-in knowledge: lay-in knowledge is substantially broader than working knowledge and, for our task, is nearly complete, containing roughly as much useful information as externally retrieved knowledge.
+Limited understanding is mainly a failure to access relevant knowledge at the right moment, not a failure to store it.
+We then examine whether models differ in their ability to construct task-specific “databases” of usable cues.
+Smaller models tend to generate simpler, more broadly transferable cues that aid retrieval and downstream use, whereas larger models generate richer but more brittle and less generalizable cues.
+Together, these results point to a complementary strategy across scales: small models propose the scaffold, and large models do the heavy reasoning.
+Contributions
+(i) We introduce a scalable task and dataset to probe LLMs’ ability to deploy domain knowledge by operationalizing a core cognitive criterion of understanding - differentiation.
+We introduce a novel large-scale dataset of hard-to-separate USPTO patents (1.3M+ information/computer and 177K+ biomedical patents granted post-2015), where surface similarity is systematically misleading and success requires technical judgment.
+(ii) We introduce a diagnostic framework that separates missing from unused, "lay-in", knowledge.
+Applying this framework reveals a deployment bottleneck: models often have relevant knowledge but fail to activate it.
+2 Hard-to-Distinguish Patents
+The dataset consists of 1,319,184 information and computer technology and 177,903 biomedical patents.
+The designed task is less likely to be impacted by pretraining as shown in Appendix G.
+Several features make technology patents particularly well-suited for our test:
+First, patents describe components of larger systems and thus occur in dense clusters of near-duplicates that are highly similar in structure and language but differ in a small number of substantively meaningful dimensions (e.g., constraints, use cases, or technical improvements) (Ouellette, 2011).
+These fine-grained distinctions - rigorously evaluated by human judges (see the third point) - directly test whether LLMs can move beyond surface similarity to recognize functional differences.
+We illustrate this pattern in Figure 1(b), which shows the distribution of similarities between one patent and its most similar – but non-identical – counterpart based on the textual embedding from (Ghosh et al., 2024), pretrained on patent-citation relations.
+The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
+We collect these similar but different patents and test whether LLMs can tell the nuances.
 In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention.
 We find that the differentiation task requires genuine conceptual linking and functional comparison, making it a strong testbed for evaluating LLM understanding in the technology domain.
-Several features make technology patents particularly well-suited for testing fine-grained semantic distinctions:
 (a) Patents are hard to read and understand
-(b) Patents are hard to distinguish
-Figure 1: Panel (a): Model perplexity of LLaMA-3.1-8B-Instruct on paper and patents abstracts.
-Panel (b): Histogram of all 1,319,184 patents, showing the distribution of each patent’s maximum cosine similarity to another (distinct) patent.
-First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement (Trippe, 2015).
-These subtle distinctions have been rigorously evaluated by human judges (we will develop this in the third point) and represent precisely the type of contrasts that test whether an LLM can move beyond surface-level similarity and recognize functional difference.
-We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from (Ghosh et al., 2024), pretrained on patent-citation relations.
-The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
-Second, patent documents are often strategically obfuscated, written to disclose the minimum required for legal protection while revealing as little as possible about the underlying invention.
-As a result, their language tends to be intentionally dense, formal, and legally contorted.
-Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms (Kong et al., 2023; Ouellette, 2012).
-This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper abstracts from ArXiv and 4,000 matched patent abstracts from the same years.
+(b) Patents and counterparts are hard to distinguish
+Figure 1: Panel (a): Model perplexity of LLaMA-3.1-8B-Instruct on paper texts and patents descriptions.
+The texts of patents are much more confusing to LLMs compared to scientific papers.
+Panel (b): Histogram of the distribution of each patent’s maximum cosine similarity to another (distinct) patent.
+Second, patent documents are often strategically obfuscated, disclosing only what is legally required while minimizing substantive revelation.
+Consequently, their language is deliberately dense, formal, and legalistic.
+Readability studies show that patent texts demand substantially higher comprehension than comparably technical scientific prose due to jargon-heavy and legally constrained drafting conventions (Kong et al., 2023; Ouellette, 2011).
+This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper texts from ArXiv and 4,000 matched patent descriptions from the same years.
 Perplexity, while not a direct measure of understanding, reflects the model’s inability to predict the next word in a sequence.
-Consistently higher perplexity on patent abstracts suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend.
-An example of paired patent abstracts and a related paper abstract can be seen in Appendix D.
-Third, and most importantly, even though patents may seem similar, we know they are not.
-Each granted patent in our dataset has undergone an extensive expert evaluation process conducted by patent examiners at the USPTO.
-This review assesses whether the application satisfies the legal requirements of novelty, non-obviousness, and utility.
-During this process, examiners conduct detailed prior art searches, compare the proposed invention to existing patents and publications, and determine whether the claimed invention constitutes a meaningful and non-trivial contribution over what already exists.
-This institutional validation is crucial for our task: it provides a strong, domain-informed guarantee that two different granted patents indeed describe conceptually distinct inventions.
+Consistently higher perplexity on patents suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend.
+An example of paired patents and a related paper can be seen in Appendix D.
+Third, and most importantly, despite surface similarity, granted patents are known to be conceptually distinct because each has undergone extensive expert review by USPTO examiners.
+This process evaluates novelty, non-obviousness, and utility through detailed prior-art searches and comparative analysis.
+As a result, institutional validation provides a strong, domain-informed, human-expert-evaluated guarantee that different granted patents represent non-trivial conceptual distinctions, even when they appear nearly identical.
 3 Evaluating LLMs’ Understanding of the Technology Domain
-We now describe how we operationalize the pairwise differentiation task.
-We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER (Ghosh et al., 2024), a contrastive learning model trained on patent citation relationships.
-For each patent, we select its most similar—but non-identical—counterpart based on cosine similarity in the embedding space.
-These pairs are then presented to the LLM, which is asked whether the two abstracts describe the same invention.
+Hard-to-distinguish patent pairs (two patents show high text similarity but with different granted IDs) are presented to the LLM, which is asked whether the two texts describe the same invention.
 [^3] See Appendix E for full prompts.
 For each prompt, the model returns two outputs: a binary label indicating “yes” (1) or “no” (0), and a confidence score on a scale from 0 to 10.
-To construct our main measure of understanding, we aggregate predictions using a confidence-weighted voting scheme applied over three independent generations of the same prompt.
+Note that the expert-evaluated ground truth is always "they are different patents".
+To construct our main measure of understanding, we aggregate LLM’s output using a confidence-weighted voting scheme applied over three independent generations of the same prompt.
 If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0.
-This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting (Geng et al., 2024).
+This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting (Geng et al., 2024; Taubenfeld et al., 2025).
 In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases.
 This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain (Geng et al., 2024).
-Alongside final labels, we also track changes in confidence and measure self-consistency, defined as the proportion of times all three responses yield the same answer.
-Lower consistency indicates greater uncertainty in model judgment.
-Figure 2: Ask LLMs "Do you think they are the same patent/different patents?", given that they are actually different. All models are Llama 3.1 instruction-tuning versions.
-Figure 2 illustrates how models of different sizes perform on the differentiation task within the technology domain.
-The first panel shows that smaller models, such as LLaMA-8B, struggle to distinguish between semantically similar but distinct patents, classifying over 70% of patent pairs as the same.
-In contrast, the larger LLaMA-405B model labels only around 10% of these pairs as identical, indicating a stronger ability to detect subtle distinctions.
-The second and third panels show that the smaller model is also less confident and consistent in its responses (across blue bars).
-The 8b model’s confidence and self-consistency of wrong answers (blue) is way higher than those of the right answer (orange), again suggesting higher uncertainty and less stable internal representations compared to the larger model.
-The fourth panel evaluates this phenomenon directly by measuring whether the model gives consistent answers across two question forms.
-For each patent pair, we ask both “Do you think these patents describe the same invention?” and “Do you think they describe different inventions?” A model with a stable internal representation of the concepts should give logically consistent answers across these two formulations, measured by joint-correct rate, the proportion of pairs where the model gives both correct answers.
-Deviations from this indicate sensitivity to surface-level wording.
-The results again show that smaller models are less consistent and more affected by such superficial cues in the phrasing, reinforcing the idea that they lack a robust conceptual understanding of the task.
-In the appendix F.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed.
-All models are more likely to judge the patents as different when asked explicitly whether they are different, even though the content of the question is logically equivalent.
-We also report additional results for other model families with varying sizes, including Qwen, Mixtral, and GPT-4o-mini.
-Again, we find that larger models perform better on the differentiation task.
-Interestingly, we observe consistent response biases in some models: Mixtral models tend to answer “yes” regardless of the question’s phrasing, while Qwen models tend to answer “no.” Based on these findings, we select the LLaMA-3.1-8B Instruct model as the primary focus for our subsequent experiments and use the question “Do you think they are the same patent?” as the default formulation for the pairwise differentiation task.
-4 Self-questioning LLMs for Improved Technology Understanding
-We begin by examining whether prompting models to ask themselves structured questions can enhance performance and, if so, how.
-We design two types of self-generated questions.
+The result of confidence-weighted voting correlates with majority voting without any confidence signals, and the correlation is 0.95.
+High confidence also well correlates with high judge accuracy, with a correlation of 0.92.
+Results using the majority voting can be found in Appendix F.
+We also track changes in confidence.
+Throughout the experiments we use 5000 randomly sampled pairs to approximate the result on the full dataset as they are statistically sufficient.
+We find results on 1000, 5000, and 10000 randomly sampled pairs show a high correlation of over 0.98.
+Figure 2: Blue bars always represent the answer "yes", orange bars always represent "no".
+Top row: Ask LLMs "Do you think they are different patents?", given that they are actually different.
+Blue: "yes" (they are different).
+Orange: "no" (they are the same).
+Bottom row: "Do you think they are the same patent?", given that they are actually different.
+Blue: "yes" (they are the same).
+Orange: "no" (they are different).
+Llama models are instruction-tuning versions.
+Looking at the top row of Figure 2, we evaluate LLM understanding by presenting two close-but-distinct patents and first asking <whether they are different?> Blue bars here always represent the answer "yes", and in this case, the proportion of responses "yes, they are different" (i.e., correct answer), along with their associated confidence, while the orange bars always denote the answer "no", and in this case, the responses "no, they are not different" (i.e., they are the same, incorrect answer).
+These results suggest that most models - with the exception of Qwen 2.5 7B - perform remarkably well; Mistral 7B, Llama 3.1 8B, and larger models like Llama 3.3 70B and GPT-4o mini all appear to identify the patents as distinct at a high rate (left panel, top row).
+However, further investigation reveals that this seemingly high accuracy is superficial rather than a reflection of true understanding.
+We introduced a logically consistent inverse question in the bottom row of Figure 2: <Are these patents the same?> For a model to demonstrate a robust grasp of the nuances between these hard-to-distinguish patents, it should answer "yes" to the first prompt (are they different?) and "no" to the second (are they the same?).
+Interestingly, our results reveal significant response biases that undermine the top-row results.
+For instance, the Mistral model tends to answer "yes" regardless of the prompt’s framing, whereas Qwen 2.5 7B shows a persistent tendency to answer "no".
+Furthermore, we observe a paradoxical "overconfidence" effect, where small models (Qwen 2.5 7B, Mistral 7B, and Llama 3.1 8B) always exhibit higher confidence when they answer "yes", and for larger models like Llama 3.3 70b and GPT-4o-mini, wrong answers show a higher confidence than correct ones.
+We also find GPT-4o mini demonstrates the strongest capability, followed by Llama 3.3 70B and then Llama 3.1 8B.
+This is evidenced by the fact that while all three models frequently identify that the patents are different ("yes" responses when asked if the patents are different) in the top row, GPT-4o mini maintains the lowest—and therefore most accurate—rate of "yes" responses when asked if the patents are the same in the bottom row.
+This suggests that as model scale and sophistication increase, so does the accuracy and the ability to maintain logical consistency when distinguishing between intricate technical documents.
+4 Experimental setups
+For failures in knowledge-intensive tasks, the most intuitive solution is to supplement the model’s knowledge, with the most direct method being Retrieval-Augmented Generation (RAG).
+Based on this premise, we designed a set of control experiments to analyze why models fail.
+Our motivation is straightforward: we allow the model to self-diagnose by asking, "What specific background knowledge would help you better distinguish between these two patents?"  [^4] See Appendix E for full prompts..
+We then tested two distinct conditions: (1) Scientific RAG: given the set of generated questions, supplying external scientific knowledge to address missing information, if any.
+(2) Self-QA: given the same set of generated questions, LLMs answer their own questions, testing whether the model can bridge the gap using its own internal knowledge, indicating a utilization failure.
+We then compare the content of these two types of answers and their impact on improving the judgment to determine the major drivers of failures.
+Specifically, we design two types of self-generated questions.
 The first, surface-level questions, target conceptual recall (e.g., What is X?
 What does Y do?), while the second encourages deeper understanding through questions about the integration of concepts, improvement, and the intent underlying invention.
-We construct the pipeline illustrated in Figure 3.
 When the model is exposed to a pair of closely related patents, $P_{1}$ and $P_{2}$, it generates six questions $Q_{ij}$ for each, where $i\in{1,2}$ indexes the patent and $j\in{1,\dots,6}$ indexes the question.
 Three of these are surface-level questions, and three target deeper comprehension.
-These questions guide the retrieval of relevant scientific knowledge, mimicking how researchers ask clarifying questions to deepen understanding.
+Randomizing the order of questions does not impact the result much.
+The prompt for generating questions could be found in Appendix E.
+In the Scientific-RAG setting, these questions guide the retrieval of relevant scientific knowledge.
 For external knowledge, we use parsed, pre-processed full texts from arXiv computer science papers.
@@
 These are then passed to the LLM as external knowledge $Knowl_{ij}$.
-The model answers its own questions using both its internal knowledge and the retrieved content, producing $A_{ij}=f_{LLM}(Q_{ij}\mid Knowl_{ij})$.
-Then the LLM makes a decision by $Decision=Judge(Patent_{1},Patent_{2}\mid\{(Q_{1j},A_{1j})\}_{j=1}^{6},\{(Q_{2j},A_{2j})\}_{j=1}^{6})$ with the previously mentioned confidence-weighted voting schema.
-Figure 3: Self-questioning LLMs: The model probes its internals to generate questions, retrieving scientific knowledge or drawing from its own knowledge base to answer them (left).
-It then uses these structured question-answer pairs to revisit the task under the confidence-weighted voting schema (right).
+The model answers its own questions using the retrieved content $Knowl_{ij}$, producing scientific-augmented answers, and we feed the six scientific QA pairs for each patent to LLMs, plus the paired patent information, to let them judge again whether the two patents are the same.
+In the Self-QA setting, all experimental conditions remain the same as above, except that no external scientific retrieval is used.
+Instead, the LLM generates its own answers and then conditions its judgment on these self-generated QA pairs together with the patent information.
 5 Results
-In this section we decompose the error rate into its components - how much of the error is due to missing information and how much is due to the inability to retrieve self-knowledge.
+In this section we decompose the error rate into its components - how much of the error is due to missing information and how much is due to the inability to deploy self-knowledge.
 Section 5.1 shows our main result.
 In the following section, we analyze how model size affects question quality and judgment accuracy (5.2).
-5.1 LLMs Do Not Know What They Know
-(a) LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA.
+Across the experiments below, we ask “do you think they are the same patent” as shown in Figure 2.
+All statistical significance analyses use the t-test.
+Ablation studies on the generated questions and retrieval quality can be found in Appendix I.
+5.1 Missing vs. Unused Knowledge
+(a) Self-talk helps improve judgment, but not as much as scientific QA.
+(b) Scientific retrieval does not introduce more “content-level” new information.
+Figure 3: Panel a: The contribution to understanding decreases in the following order: scientific QA, self-QA, and question-only.
 $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs.
 scientific answers).
-(b) Scientific retrieval does not introduce more “content-level” new information for LLM to answer these questions.
-Figure 4: Mechanisms of "LLMs do not know what they know".
-Figure 4(a) presents our initial results from a random sample of 1000 patent pairs, examining how questions and answers affect classification performance.
-The baseline represents the initial accuracy of Llama 8B without any self-questioning or scientific retrieval.
+Panel b: no significant difference could be found between scientific information and self-produced answers.
+Figure 3(a) presents our results from the random sample of 5000 patent pairs, examining how questions and answers affect classification performance.
+The baseline represents the initial accuracy of Llama 3.1 8B without any self-questioning or scientific retrieval.
 At this baseline level, Llama 8B incorrectly classifies approximately 70% of patent pairs as identical.
+Increasing the number of question–answer pairs leads to a more robust understanding.
 The blue line demonstrates the impact of adding questions alone.
-This modest improvement suggests that simply posing understanding questions helps the model perform better by providing structure that aids text parsing, though the effect of merely enumerating questions remains limited.
+This modest improvement suggests that simply posing understanding questions surprisingly helps the model perform better by providing structures and cues that aid text parsing and internal retrieval, though the effect of merely enumerating questions remains limited.
 The orange line reveals that adding self-generated answers significantly improves performance.
-This substantial gain implies that the model already contains useful knowledge for the differentiation task, but this knowledge is not being effectively utilized.
-These results suggest that LLMs do not know what they know—bringing forth relevant information that enhances understanding can dramatically improve performance.
-Finally, the green line shows the improvement achieved by providing retrieved answers.
-This additional performance gain indicates that while the model’s internal knowledge is valuable, it does not match the quality of complete information available in the training data.
-Notably, with six question-answer pairs, performance reaches the same level as the 405B parameter model, effectively eliminating the error rate.
-Importantly, most of the error reduction is achieved through self-generated answers, suggesting that the model contains the most useful information within its internal knowledge representation.
+This substantial gain implies that the model already contains useful knowledge for the differentiation task, but this knowledge is not being effectively utilized unless explicitly prompted.
+The green line shows the improvement achieved by providing scientifically retrieved answers.
+Importantly, error reduction is achieved more substantively through baseline to self-generated answers than self-generated answers to scientific ones, suggesting that the model contains the most useful information within its internal knowledge representation.
 However, the model lacks the ability to effectively access and apply this knowledge when tackling tasks that require such domain expertise.
-Our results demonstrate that models fail to retrieve the information necessary for the task, and the information they do retrieve is of lower quality than that obtained from additional sources.
-We next explore the factors driving this performance difference, considering two possible explanations.
-First, LLMs function as compressed representations of knowledge but do not preserve complete information.
-Scientific and technological information tends to be sparsely and unevenly distributed in training corpora, and the compression process significantly degrades the fidelity of such underrepresented knowledge.
-Second, LLMs may acquire new knowledge at inference time—meaning the external knowledge we provide is entirely unknown to the LLM and must be learned and incorporated during inference.
+We next explore the factors driving this performance difference between self-produced answers vs scientific answers, considering two possible explanations.
+First, LLMs function as compressed representations of knowledge but do not preserve the richness of information.
+Scientific and technological information tends to be sparsely and unevenly distributed in training corpora, and the compression process significantly degrades the richness of such underrepresented knowledge (Jiang et al., 2020; Jaiswal et al., 2024; Hoang et al., 2023; Sun et al., 2024).
+This implies that the “missing information” hypothesis largely does not hold.
+The LLMs appear to internally store the relevant information; however, they either fail to effectively deploy it or suffer from a loss of informational richness due to compression.
+We thus also observe that cues only (questions) still help the deployment and dynamic use process within the compressed knowledge store.
+Second, LLMs may acquire new knowledge during external retrieval - meaning the external knowledge we provide is entirely unknown to the LLM and must be learned and incorporated during judgment.
+That implies that the “missing information” is the key.
 We argue that the first explanation is likely the primary cause of the improvement we observe.
-The models have likely already encountered the data we use: Llama 8B was released in July 2024444https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the retrieved papers that generated scientific answers had already been published.
-However, while the models may have seen these papers during training, this exposure might not have significantly influenced the learned weights.
-To further investigate what type of information is contained in inference-time retrieval versus the model’s internal knowledge, we analyze the answers by embedding all responses from both scientific sources and the LLM’s self-generated answers using the model from (Ghosh et al., 2024).
+The models have likely already encountered the scientific data we use as external sources: LLaMA 3.2 was released around October 2024555https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, when approximately 86% of the retrieved papers that generated scientific answers had already been published.
+To further investigate what type of information is contained in inference-time retrieval versus the model’s internal knowledge, we analyze the answers by embedding all responses from both scientific sources and the LLM’s self-generated answers using the model from (Ghosh et al., 2024), specifically pretrained on patent citations.
 To visualize and compare the distribution of embedding spaces, we reduce dimensionality to 2D and estimate boundaries using 5% Gaussian kernel density estimation—drawing contours that enclose the densest 95% of data points in each embedding set.
-If LLMs substantially incorporate new scientific content during inference, we would expect their embedding regions to diverge meaningfully from those derived from self-generated responses.
+If LLMs substantially incorporate new scientific content for answers by external retrieval, we would expect their embedding regions to diverge meaningfully from those derived from self-generated responses.
 However, as shown in Figure 3 panel b, the embedding regions largely overlap.
 Notably, self-generated responses display a distinct outlier region, which we suspect reflects hallucinated content introduced by the LLM.
-Despite occupying similar regions in the 2D embedding space that captures the greatest variance in the high-dimensional (768D) space, these two answer types differ in several meaningful ways.
-Scientific answers average $62.1\pm 1.8$ words and are presumably richer in detail and informativeness, drawing more heavily from scientific evidence.
+We also qualitatively interpret these questions and answers as shown in Appendix H.
+Despite occupying similar regions, these two answer types differ in several meaningful ways.
+Scientific answers average $62.1\pm 1.8$ words and are presumably richer in detail and informativeness.
 In contrast, self-generated answers are notably shorter, averaging only $38.2\pm 0.6$ words—approximately half the length of scientific responses.
-We therefore, interpert the results as suggesting that although the LLMs already have the knowledge needed to answer the question, and obtain better performances - this knowledge is compressed, less detailed and not as rich as the original scientific data.
+We therefore, interpret the results as the following insights.
+The issue of compression Although in our settings LLMs largely possess the knowledge required for knowledge-intensive tasks, this knowledge is highly compressed, less detailed, and substantially less rich than the original scientific data from which it was learned.
+The issue of deployment Unless explicitly activated through mechanisms such as self-generated answers or structured elicitation cues (e.g., questions alone), the compressed knowledge in LLMs remains difficult to access and largely unavailable for dynamic deployment.
+This perspective also challenges an assumption of RAG: that supplying retrieved documents can remedy knowledge deficiencies at inference time.
+If performance gaps arise from limitations in knowledge deployment or integration - rather than simple absence - retrieval may be insufficient.
+Retrieved content can remain shallowly integrated or misaligned with internal representations.
+Thus, RAG treats knowledge access as the bottleneck (Zhou et al., 2026; Cheng et al., 2025), whereas our findings suggest that knowledge representation and dynamic use are equally, if not more, critical.
+Improving performance may therefore require advances in how models internalize retrieved information, not merely expanding input.
 5.2 It’s Not Only What to Ask, but Who Asks It
-Self-questioning helps because it aligns the model’s attention with task-relevant concepts.
-This suggests a transferring question: are the most useful questions those that the same model would ask itself, or can questions authored by a different model—smaller or larger—better align the answerer’s knowledge with the task?
+Figure 4: Question models: llama 1b, 8b, 70b, 405b.
+Answer models: 3b and 8b.
+Panels a and b: Self-generated questions are compatible with the model’s own answering capacity.
+Questions generated by smaller models are compatible with larger answer models but not the converse.
+"Compatible" means a lower misjudgment rate (judging as "same patents").
+Panels c to h: mechanisms.
+Panels c and d: higher perplexity when answering small models’ questions.
+Panels e and f: lower similarity between [small models’ questions + answer models’ answers] and [the patent texts], implying searching for more new information when answering small models’ questions.
+Panels g and h: smaller models’ questions are more open-ended (the SD of embedding distances of each answer from the mean of all answers).
+A transferring question emerges: are the most useful questions those that the same model would ask itself, or can questions authored by a different model—smaller or larger—better align the answerer’s knowledge with the task?
 We next test how question–answer sets travel across scale and where misalignment helps or hurts.
@@
 The 3B and 8B models provide a more informative middle ground: they show some understanding but still benefit from superior prompts.
-As shown in Figure 5 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B.
+As shown in Figure 4 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B (i.e., misjudge rate (judging as "same patents") is lower than questions from bigger models).
 This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion.
@@
 This may be because larger models can readily interpret, build on, and give a better answer to smaller models’ questions.
-Large model questions appear to go beyond the bounds (and dimensionality) of smaller model understanding and answering capacity.
-Figure 5: Panels a and b: Questions generated by smaller models are compatible with larger models, but not the converse.
-Panels c to h: mechanisms.
-Panels c and d: higher perplexity when answering small models’ questions.
-Panels e and f: more new information when answering small models’ questions.
-Panels g and h: smaller models’ questions are more open-ended.
+Large model questions appear to go beyond the bounds (and dimensionality) of smaller models’ understanding and answering capacity.
 To probe this further, we generate 10,000 question-scientific answer pairs using the 3B and 8B models (the answer model), based on questions from either smaller or larger models relative to the answer models’ own size (smaller or bigger question models), resulting in 40,000 total pairs.
-We then measure textual similarity between each QA pair and its associated patent abstract.
-When a larger model answers questions from a smaller model, the resulting QA content is consistently less similar to the original patent abstract, suggesting that small models’ questions bring in more external scientific information to improve the quality, texture, and richness of knowledge beyond the raw patent texts (Figure 5 panels e and f).
+We then measure textual similarity between each QA pair and its associated patent text.
+When a larger model answers questions from a smaller model, the resulting QA content is consistently less similar to the original patent text, suggesting that small models’ questions bring in more external scientific information to improve the quality, texture, and richness of knowledge beyond the raw patent texts (Figure 4 panels e and f).
 Small model questions “discover more” external information from the sources they search.
 We also re-feed the scientific QA pairs into the 3B and 8B models and record their model perplexity.
-This lower similarity predictably activates more latent reasoning in the models, as reflected in their answer’s higher perplexity (Figure 5 panels c and d).
+This lower similarity predictably activates more latent reasoning in the answering models, as reflected in their answer’s higher perplexity (Figure 4 panels c and d).
 This may stem from how smaller models phrase their questions.
 With limited capacity, their questions tend to have two important features: (1) Small model questions are more fundamental — 68% of 1B’s questions contain “what is” or “what are” (usage/purpose of components, concepts, and so on), compared to only 44% for 405B.
+They are precisely the fundamentals of judging patents.
 (2) Smaller models tend to ask shorter, less contrastive, and more open-ended questions, potentially yielding higher information entropy (Figure 4 panels g and h).
@@
 We quantify “open-endedness” as the embedding distance of each answer from the mean of all answers (standard deviation).
-Questions from smaller question models elicit more variable answers.
+Questions from smaller question models elicit more varied answers.
 6 Final Remarks
-Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge.
-While models hold useful internal insights, these are often latent and underused.
-Self-questioning helps activate this hidden compressed capacity, but falls short for complex tasks.
-Interestingly, the fundamental questions from smaller models often align with the needs of larger models.
-This suggests a scalable strategy: use smaller models to structure problems and guide deeper reasoning in stronger models.
-Impact: our work and the associated new benchmark on patent distinction offer broad insights into how AI can be used to automate the evaluation of scientific and technical contributions.
+We reframe LLM failures not as gaps in knowledge acquisition, but as limitations in knowledge deployment.
+Using a patent differentiation task that requires the dynamic use of knowledge, we show that models often possess relevant knowledge yet fail to activate it, whereas external retrieval yields diminishing returns when utilization—not access—is the primary bottleneck.
+Complementarities across model scales further suggest hybrid strategies that combine structural scaffolding from smaller models with deeper reasoning from larger ones.
+This work brings insights into the right way of building science-powered LLMs for automating the assessment of innovation.
+Limitations
+Although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself.
+Rather than relying on fixed sources or searching questions from scientific databases, LLMs could learn to refine or even generate better retrieval queries via reinforcement learning, active learning, or model-in-the-loop optimization based on science or what they have learned from prior reasoning steps.
+An agentic framing - where models, like a baby sensing the world, autonomously refine their questions and knowledge, decide whether to retrieve information, evaluate whether to accept or reject what is retrieved, and determine how much knowledge to seek for one question/how many questions to ask - could enable deeper forms of exploration beyond single-pass self-questioning.
+The exact nature of “resolution” in distinguishing patents remains underdefined.
+That is, even when models succeed at effectively splitting, it is not always clear what semantic threshold they are using.
+Are they resolving based on functionality, application, architecture, or terminology?
+Future work could explicitly model and evaluate these questions to understand what kinds of conceptual distinctions LLMs are truly capable of making by probing their internal, layer-wise conceptual representations and dynamics.
+It will also be interesting to look into differences in judgment performance across patent classes and categories, as LLM internal knowledge likely varies across fields.
+Ethics Statement
+This work complies with the Code of Ethics.
+We used AI assistants to expedite the coding process.
+All AI-generated code snippets were carefully reviewed and verified by the first author prior to incorporation.
+For manuscript preparation, AI assistants were used exclusively for grammar checking.
+No external ethics review was required, as the experiments involved only the analysis of patent texts and therefore posed no risk to individuals across any demographic or geographic groups.
+The patent data are licensed under the Attribution 4.0 International license (see https://patentsview.org/rules-of-conduct?utm_source=chatgpt.com), which permits copying and redistribution in any medium or format for any purpose, including commercial use.
+The data are publicly available through the USPTO; our contribution was limited to collection and cleaning.
+All experiments were conducted using two NVIDIA A100 GPUs with 40 GB of memory each.
 Appendix
 Appendix A Understanding Through Differentiation
-Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; Grimm [2024] it fundamentally involves grasping the relationships that connect concepts and principles within a domain.
+Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures.
+It fundamentally involves grasping the relationships that connect concepts and principles within a domain.
 This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall.
 Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
-Cognitive science, from [Gibson, 1969] through [Tversky, 1977] and Gentner’s structure-mapping work [Markman and Gentner, 2000] to recent relational-reasoning research [Alexander and Dumas, 2018], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”.
+Cognitive science, from (Gibson, 1969) through (Tversky, 1977) and Gentner’s structure-mapping work (Gentner and Markman, 1997) to recent relational-reasoning research (Alexander et al., 2016), demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”.
 Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’.
@@
 Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes BB.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims.
-Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [Pearl, 2009, Lake et al., 2017, Heckman and Pinto, 2022].
-Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [Gopnik and Wellman, 2012].
-LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language Wang et al.
-[2018], Wang et al.
-[2019], OpenAI [2023], Brown et al.
-[2020].
-Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model Vafa et al.
-[2024] that describes and explains how things relate to one another.
+Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation (Pearl, 2009; Lake et al., 2017; Heckman and Pinto, 2022).
+Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes (Gopnik and Wellman, 2012).
+LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language (Wang et al., 2018), (Wang et al., 2019), (OpenAI, 2023), (Brown et al., 2020).
+Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model (Vafa et al., 2024) that describes and explains how things relate to one another.
 The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts.
 Consequently, measuring an LLM’s proficiency in discerning similarity and difference within highly specific concepts and fields offers a powerful proxy for evaluating its understanding of the conceptual foundations and landscape of these fields.
-Our measure of understanding draws inspiration from previous benchmarks used to assess language comprehension in earlier generations of language models.
-GLUE Wang et al.
-[2018], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other.
-Similarly, our tasks focus on whether models can distinguish subtle conceptual differences.
-While prior benchmarks often relied on synthetic or crowd-sourced data, e.g.
-Bowman et al.
-[2015], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring a fine-grained understanding of technical content.
 Appendix B Related Literature
-Relations to other benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding.
-But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [McIntosh et al., 2024, Wu et al., 2024, Davis, 2023].
+Relations to current benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding.
+But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023).
 High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension.
@@
 Even when test content is novel, models can exploit formatting quirks.
+Current benchmark agreement tests are strongly influenced by benchmark-orthogonal factors such as question format (Wu et al., 2025).
 Alzahrani et al.
-[Alzahrani et al., 2024] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions.
+(Alzahrani et al., 2024) showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions.
 Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application.
-Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [Arakelyan et al., 2024].
-Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [Burnell et al., 2023].
+Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics (Arakelyan et al., 2024).
+Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities (Burnell et al., 2023).
 In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities.
-Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [Asthana et al., 2024] or contrast and counterfactual sets [Lewis and Mitchell, 2024]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [West et al., 2024]..
+Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase (Asthana et al., 2024) or contrast and counterfactual sets (Lewis and Mitchell, 2024)  [^7] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content (West et al., 2024)..
 Organizing internal and external knowledge: A core challenge in LLM understanding is how to activate the vast latent knowledge embedded in the model’s billions of parameters.
-Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [Borgeaud et al., 2022, Hoang et al., , Hoang et al., 2023, Jaiswal et al., 2024, Pan et al., 2025].
+Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand (Borgeaud et al., 2022; Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024; Pan et al., 2025).
 Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted.
 The model might skillfully "reason" with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge.
-This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [Schulhoff et al., 2024, Gopnik and Wellman, 2012].
+This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial (Schulhoff et al., 2024; Gopnik and Wellman, 2012).
 To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources.
-Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [Prince et al., 2024].
+Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence (Prince et al., 2024).
 Lewis et al.
-[Lewis et al., 2020] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks.
+(Lewis et al., 2020) demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks.
 Borgeaud et al.
-[Borgeaud et al., 2022] introduced the RETRO (Retrieval-Enhanced Transformer) model.
+(Borgeaud et al., 2022) introduced the RETRO (Retrieval-Enhanced Transformer) model.
 A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval.
 He et al.
-[He et al., 2022] showed that an LLM with a retrieval step produces more faithful, grounded understanding – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness.
+(He et al., 2022) showed that an LLM with a retrieval step produces more faithful, grounded understanding – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness.
 In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries.
-This is likely because LLMs gain new information during retrieval.
-LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [Shumailov et al., 2024].
+LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data (Shumailov et al., 2024).
 By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarities and differences.
-Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [He et al., 2020, Gao et al., 2021, Chen et al., 2020, Cohan et al., 2020].
+Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity (He et al., 2020; Gao et al., 2021; Chen et al., 2020; Cohan et al., 2020).
 Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure.
 While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.
-Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [Huang et al., 2023, Qu et al., 2024, Song et al., 2024, Wang et al., 2023c, b].
-Instead of only generating direct answers, many contemporary models are encouraged to ask themselves clarifying questions and work through sub-problems, mimicking a Socratic or pedagogical approach.
-Early prompting techniques like chain-of-thought (CoT) [Wei et al., 2022] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks.
-Building on this, researchers introduced methods for the model to generate its own intermediate questions.
-For example, self-ask prompting [Press et al., 2022] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces.
-Beyond decomposition, recursive self-reflection techniques have emerged to let LLMs iteratively improve their answers.
-Madaan et al.
-[Madaan et al., 2023] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds.
-Shinn et al.
-[Shinn et al., 2023] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts.
-Huang et al.
-[Huang et al., 2023] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks.
-Miao et al.
-present SelfCheck [Miao et al., 2024], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks.
-Similarly, techniques like least-to-most prompting [Zhou et al., 2022] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges.
-Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [Wang et al., 2023a]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work.
-These introspective or pedagogically-inspired strategies are motivated by human learning: students often learn better by asking themselves questions and reflecting on mistakes.
-Empirically, such approaches can bolster an LLM’s conceptual grasp and reduce reasoning errors.
-Nevertheless, there are also limitations and mixed findings.
-It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [Balepur et al., 2025].
-As we demonstrate, large models are much more likely to productively answer small model questions than the converse.
-Not all tasks benefit uniformly from self-reflection.
-Li et al.
-[Li et al., 2024] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task.
-Thus, while self-questioning and introspective prompts represent a powerful paradigm for enhancing LLM reasoning, careful tuning is needed to ensure the model’s “inner dialogue” remains productive and does not overthink or hallucinate.
 Appendix C Judging Close Patents VS. Rewritten Patents
 (a) Few-shot CoT fails to help distinguish between patents with meaningful technical differences.
-(b) Few-shot CoT helps identify similarities across patent abstracts and their rewritten versions.
-Figure 6: Few-shot CoT reasoning can distinguish the rephrased abstracts (rewritten by the GPT-o3 model), but cannot distinguish their counterpart (another unique patent).
-Following on prior sections that focus on distinguishing deceptively similar patents, in this section, we explore whether and how LLMs are able to merge identical but paraphrased patent ideas.
-Our approach to measuring LLMs’ ability to generalize and identify the same ideas across distinct articulations is to generate rephrased versions of each patent abstract using the OpenAI ChatGPT o3 model with the prompt “rephrase the patent, but keep its meaning.” We then task the LLaMA-3.1-8B-instruct model with identifying whether a patent abstract and its paraphrase describe the same invention.
-This was designed to parallel the original distinguishing experiment, in which we randomly sampled 1,000 patents and identified pairs that were most semantically similar.
-We then asked the LLaMA-3.1-8B-instruct model to determine whether the two abstracts described the same invention, using a human-annotated ground truth indicating that they were different.
+(b) Few-shot CoT helps identify similarities across patent texts and their rewritten versions.
+Figure 5: Few-shot CoT reasoning can distinguish the rephrased patent texts (rewritten by the GPT-o3 model), as shown in the right panel (lower misjudge rate), but cannot distinguish their counterpart (another unique patent, as shown in the left panel).
+Y-axis always represents the response of "yes", i.e., the error rate (Left: yes when asking "are they the same" for close-but-different patent pairs.
+Right: yes when asking "are they different" for rephrased patents.)
+Following on prior sections that focus on distinguishing deceptively similar patents (the splitting task), in this section, we explore whether and how LLMs are able to merge identical but paraphrased patent ideas (the merging task).
+Our approach to measuring LLMs’ ability to generalize and identify the same ideas across distinct articulations is to generate rephrased versions of each patent text using the OpenAI ChatGPT o3 model with the prompt “rephrase the text, but keep its meaning.” We then task the LLaMA-3.1-8B-instruct model with identifying whether a patent text and its paraphrase describe different interventions (the correct answer: no).
 In raw (zero-shot) judgments, the model performed poorly on both merging and splitting tasks.
-We then introduced zero-shot CoT prompting (“think step-by-step”) and few-shot CoT, where we supplied examples of both merging (paraphrased) and splitting (semantically similar but distinct) patent pairs.
-Few-shot CoT substantially improved the model’s accuracy in recognizing paraphrased abstracts as equivalent—but had minimal effect on its ability to distinguish between closely related yet different inventions.
-This pattern suggests that CoT reasoning helps the model overcome basic semantic similarity checks and identify paraphrases, likely by encouraging step-by-step alignment at the surface level.
+We then introduced zero-shot Chain-of-Thought (CoT) prompting (“think step-by-step”) and few-shot CoT, where we supplied examples of both merging (paraphrased) and splitting (semantically similar but distinct) patent pairs.
+Few-shot CoT substantially improved the model’s accuracy in recognizing paraphrased texts as equivalent—but had minimal effect on its ability to distinguish between closely related yet different inventions.
+This pattern suggests that more reasoning helps the model overcome basic semantic similarity checks and identify paraphrases, likely by encouraging step-by-step alignment at the surface level.
 Our hypothesis is that the model passes the rephrasing test primarily through pattern matching and semantic overlap, rather than through true conceptual integration.
-This explains why CoT improves performance on the rephrasing task: the test is relatively easy in the sense that our rewritten version keeps semantic similarity high, such that shallow strategies suffice.
-In contrast, differentiating between semantically similar but substantively different patents remains hard, even with CoT, because it requires more than aligned phrasing—it demands conceptual depth.
-In principle, a more advanced rephrasing model—one that intentionally avoids superficial alignment and instead rewrites concepts in deeper, structure-preserving ways—might push the LLM closer to the more difficult task of distinguishing near matches.
+The merging test is relatively easy in the sense that our rewritten version keeps key technical terms unchanged but only changes how the sentence is structured, such that shallow strategies suffice.
+In contrast, differentiating between semantically similar but substantively different patents remains hard, because it requires more than aligned phrasing—it demands conceptual depth.
+In principle, a more advanced rephrasing model—one that intentionally avoids superficial alignment and instead rewrites concepts in deeper, functionally equivalent ways—might push the LLM closer to the more difficult task of distinguishing near matches.
 Designing such a rephrasing system is far from trivial.
-Thus, while our current rewrite approach does not fully obscure surface similarity, the model’s success only when the task is easy reinforces the motivation behind our benchmark: LLMs rely heavily on surface-level semantic similarity, rather than demonstrating a deeper understanding of the underlying concepts.
-In this way, our dataset represents a frontier capability for all contemporary LLMs, even though larger models outperform smaller ones.
+Thus, while our current rewrite approach does not fully obscure surface similarity, the model’s success only when the task is easy reinforces the motivation behind our task: LLMs rely heavily on surface-level semantic similarity, rather than demonstrating a deeper understanding of the underlying concepts.
 Appendix D The Patent Dataset
-The patent dataset comprises 1,319,184 patent pairs in total, drawn from five Cooperative Patent Classification (CPC) classes: H05 – Electric techniques not otherwise provided for; H04 – Electric communication technique; G06 – Computing; Calculating; Counting; H03 – Basic electronic circuitry; and G11 – Information storage.
+The computer science patent dataset comprises 1,319,184 patent pairs in total, drawn from five Cooperative Patent Classification (CPC) classes: H05 – Electric techniques not otherwise provided for; H04 – Electric communication technique; G06 – Computing; Calculating; Counting; H03 – Basic electronic circuitry; and G11 – Information storage.
 Within these, the patents are further categorized into 41 CPC subclasses: [’H05B’, ’H04L’, ’H05H’, ’H05K’, ’G06V’, ’G06N’, ’G06T’, ’H04N’, ’G06F’, ’G06Q’, ’H03M’, ’G06K’, ’H04B’, ’H04R’, ’G11B’, ’G11C’, ’H04M’, ’H03B’, ’H03F’, ’H03G’, ’H03H’, ’H03K’, ’H03L’, ’H04H’, ’H04J’, ’H04K’, ’H04W’, ’H04Q’, ’H04S’, ’H05F’, ’H05G’, ’H03C’, ’H03J’, ’G06E’, ’H03D’, ’G06G’, ’G06M’, ’G06C’, ’H05C’, ’G06J’, ’G06D’].
+The 177,903 biomedical patents are drawn from USPTO classes C01, C07–C14, and C21–C30.
 Out of all matched patent pairs in our dataset, 82.55% share the same CPC class, and 68.10% share the same, more specific CPC subclass.
@@
 We also show a scientific paper abstract to help readers perceive the distinct writing styles between patents and papers.
-Patent 1: A wireless device (18) is configured for use in a wireless communication system.
-The wireless device (18) is configured to receive, from a radio network node (12), a multicast or broadcast transmission (14) and downlink control information indicating a number of repetitions with which the multicast or broadcast transmission (14) is transmitted.
-The wireless device (18) is also configured to operate according to a negative acknowledgement only feedback scheme in which the wireless device (18) is configured to transmit a negative acknowledgement to the radio network node (12) if decoding of the multicast or broadcast transmission (14) fails using the indicated number of repetitions and to refrain from transmitting a positive acknowledgement to the radio network node (12) if decoding of the multicast or broadcast transmission (14) succeeds.
+Reference indexes are removed.
+Patent 1: A wireless device is configured for use in a wireless communication system.
+The wireless device is configured to receive, from a radio network node, a multicast or broadcast transmission and downlink control information indicating a number of repetitions with which the multicast or broadcast transmission is transmitted.
+The wireless device is also configured to operate according to a negative acknowledgement only feedback scheme in which the wireless device is configured to transmit a negative acknowledgement to the radio network node if decoding of the multicast or broadcast transmission fails using the indicated number of repetitions and to refrain from transmitting a positive acknowledgement to the radio network node if decoding of the multicast or broadcast transmission succeeds.
 Patent 2: Methods, systems, and devices for wireless communications are described that support feedback for multicast communications.
@@
 The base station may monitor for the feedback message and determine to retransmit the multicast information if the feedback message is received.
-An Arxiv paper including the answer from the sampled patent pair: The Third Generation Partnership Project (3GPP) has recently published its Release 16 that includes the first Vehicleto-Everything (V2X) standard based on the 5G New Radio (NR) air interface.
+An Arxiv paper including the answer to questions about judging the sampled patent pair: The Third Generation Partnership Project (3GPP) has recently published its Release 16 that includes the first Vehicleto-Everything (V2X) standard based on the 5G New Radio (NR) air interface.
 5G NR V2X introduces advanced functionalities on top of the 5G NR air interface to support connected and automated driving use cases with stringent requirements.
@@
 Prompt for Question Generation System Prompt: You are an expert in patent comprehension.
-Your task is to generate structured questions that assess a reader’s background knowledge necessary to understand a given patent abstract.
+Your task is to generate structured questions that assess a reader’s background knowledge necessary to understand a given patent.
 The questions should focus on foundational concepts, principles, and applications relevant to the patent’s domain without explicitly referencing the patent itself.
@@
 These questions should be scenario-based and encourage practical thinking.
-Do not directly reference the patent abstract in any question.
+Do not directly reference the patent text in any question.
 Instead, ensure all questions test knowledge that would be useful for understanding the patent without explicitly addressing its content.
 Maintain a structured and logical flow from basic recall to deeper conceptual understanding and practical application.
-Basic Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Abstract: {patent_summary} Generate a set of {qnum} questions that test a reader’s ability to recall fundamental concepts, key terms, and basic components necessary to understand the given patent abstract.
+Basic Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Text: {patent_summary} Generate a set of {qnum} questions that test a reader’s ability to recall fundamental concepts, key terms, and basic components necessary to understand the given patent.
 The questions should focus on defining terms, identifying parts, and recognizing the primary function of the described technology.
-Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }} Conceptual Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Abstract: {patent_summary} Generate a set of {qnum} questions that assess a reader’s comprehension of the given patent abstract by requiring them to explain how different elements of the invention work together.
+Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }} Conceptual Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Text: {patent_summary} Generate a set of {qnum} questions that assess a reader’s comprehension of the given patent text by requiring them to explain how different elements of the invention work together.
 The questions should encourage the reader to summarize the design, describe interactions between components, and interpret the intended improvements of the invention.
@@
 2.
-Synthesize those points with your knowledge and try your best to answer the question.
+Try your best to answer the question.
 Return your answer in the following JSON format: {{ “answer”: “your answer” //your answer to the question }}
 Prompt for the Splitting Task Judgment You are a patent judge.
-Below are two patent abstracts with extra information.
-This extra information may or may not be useful in assessing the patents.
-You may choose to use it at your discretion.
-// This is for placebo answers Your task: judge whether they describe the same patent.
-Abstract 1: {abstract1} {qa_block1} Abstract 2: {abstract2} {qa_block2} Strictly return your concise answer in the following JSON format: {{ “label”: 1 // 1 if they describe the same patent, 0 if not “confidence”: 10 // your confidence score from 0 to 10, with 10 representing the highest confidence }} // Note: // This applies to the reverse questioning scenario (the major experiments in the paper): // "Do you think they are the same patent?" — when, in fact, the patents are judged to be *different* by human experts.
+Below are two patent texts with extra information.
+Your task: judge whether they describe the same patent.
+Patent 1: {patent1} {qa_block1} Patent 2: {patent2} {qa_block2} Strictly return your concise answer in the following JSON format: {{ “label”: 1 // 1 if they describe the same patent, 0 if not “confidence”: 10 // your confidence score from 0 to 10, with 10 representing the highest confidence }} // Note: // This applies to the major experiments in the paper: "Do you think they are the same patent?" — when, in fact, the patents are judged to be *different* by human experts.
 // Invert this setup: Replace "same" with "different", and return 1 if the LLM concludes that the patents are *different*.
-This yields an alternative version of the experiment (See Section 3 Figure 2 in the paper).
+This yields an alternative version of the experiment (See Section 3 Figure 2 top row in the paper).
 Prompt for Self-talk Answering Try to answer the question. Question: {question} Return your concise answer in the following JSON format: {{ "answer": "your answer" // your answer to the question }}
 Appendix F Results for Majority Voting Schema
-F.1 Performances of other LLMs with varying sizes on splitting patents
-Figure 7: Ask LLMs "Do you think they are the same patent?", given that they are actually different. Judge using majority voting, and 1 = same. Models are instruction-tuning versions.
 Figure 6: Ask LLMs "Do you think they are the same patent?", given that they are actually different.
-Judge using majority voting, and 1 = same.
-Models except for GPT-4o-mini are instruction-tuning versions.
+The judge is using majority voting, and 1 = same (wrong answer).
+Llama models are instruction-tuning versions.
 Figure 7: Ask LLMs "Do you think they are different patents?", given that they are actually different.
-Judge using majority voting, and 1 = yes.
-Models except for GPT-4o-mini are instruction-tuning versions.
-F.2 LLMs do not know what they have known
+The judge is using majority voting, and 1 = yes (correct).
+Llama models are instruction-tuning versions.
 Figure 8: Asking Llama 3.1 8B instruct "Do you think they are the same patent?" on actually different patents (1 = wrong answer "yes").
-LLM answers using majority voting.
-It does not know what it has known.
-These results are statistically significant: Self-talk still helps improve the baseline judge but not as good as scientific self-questioning: $q_{1}$ to $q_{5}$ p=0.000; $q_{6}$ p=0.006.
-F.3 Smaller models’ questions help mid-size models more than those from larger models
+LLM judges patents using majority voting.
+These results are statistically significant: Self-talk still helps improve the baseline judge but not as good as scientific answers: $q_{1}$ to $q_{5}$ p=0.000; $q_{6}$ p=0.006.
 Figure 9: Llama 3B/8B’s judge with questions proposed by different models with different sizes.
@@
 We observe that bigger models’ questions are less useful for the answer models to judge.
-Appendix G Discussion of Limitations and Future Work
-Several limitations and directions for future work remain.
-Although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself.
-Rather than relying on fixed sources or searching questions from scientific databases, LLMs could learn to refine or even generate better retrieval queries via reinforcement learning, active learning, or model-in-the-loop optimization based on science or what they have learned from prior reasoning steps.
-An agentic framing - where models, like a baby sensing the world, autonomously refine their questions and knowledge, decide whether to retrieve information, evaluate whether to accept or reject what is retrieved, and determine how much knowledge to seek for one question/how many questions to ask - could enable deeper forms of exploration beyond single-pass self-questioning.
-The exact nature of “resolution” in distinguishing patents remains underdefined.
-That is, even when models succeed at effectively splitting, it is not always clear what semantic threshold they are using.
-Are they resolving based on functionality, application, architecture, or terminology?
-Future work could explicitly model and evaluate these questions to understand what kinds of conceptual distinctions LLMs are truly capable of making by probing their internal, layer-wise conceptual representations and dynamics.
-It will also be interesting to look into differences in judgment performance across patent classes and categories, as LLM internal knowledge likely varies across fields.
+Appendix G Pretraining Exposure
+The key to evaluating the model understanding is that the tests we created are unlikely to be in the training data.
+If they are, it might be that the models simply regrgitate case that they have already seen before and not demonstrate understanding.
+Note that we are less concerned with whether the patent texts alone are present in the pretraining data, because we deliberately construct a pairing mechanism whose likelihood of having been encountered during pretraining is extremely low.
+The patent set used in the paper contains roughly $n=1.5\times 10^{6}$ patent texts.
+Each text averages about $400$ words, or approximately $520$ byte‑pair‑encoded (BPE) tokens [^8] Rule of thumb: one English word $\approx 1.3$ BPE tokens.
+See the OpenAI Help Center article “What are tokens and how to count them?” https://help.openai.com/en/articles/4936856-what-are-tokens-and-how-to-count-them..
+The number of unordered pairs is $\tfrac{n(n-1)}{2}$, so concatenating every possible pair would require approximately $1.2\times 10^{15}\ \text{tokens}$.
+That total is multiple orders of magnitude larger than some of the largest publicly reported training corpora, even though a large share of these corpora is (unrelated) github code, books, Wikipedia, etc: GPT models were trained on $\sim 3\times 10^{11}$ tokens [^9] Brown, T.B.
+et al.(2020), “Language Models are Few‑Shot Learners.” and Llama‑70B on $\sim 2\times 10^{12}$ tokens [^10] Touvron, H.
+et al.(2023), “Llama: Open Foundation and Fine‑Tuned LLMs.”.
+The size of Redpajama English tokens is $\sim 2\times 10^{13}$  [^11] https://github.com/togethercomputer/RedPajama-Data.
+Appendix H Example Questions
+Qualitative inspection of the generated questions suggests that answering them should not require substantial novel information beyond the models’ training (echoing the missing vs.
+unused knowledge hypothesis).
+Many questions are straightforward and focus on established concepts or systems.
+For example, questions involving the TCP Port Service Multiplexer (TCPMUX) relate to long-standing infrastructures, with patents contributing only incremental advancements.
+As shown in our original paper, 68% of the Llama 1B model’s questions and 44% of the Llama 405B model’s questions include phrases like “what is” or “what are,” indicating the dominance of factual, definitional queries.
+A sample of questions can be seen in Figure 10.
+Figure 10: Example questions.
+Appendix I Ablation Studies
+Figure 11: More scientific-QA improves the judge over the placebo answer: For panel (a): $p_{q1}$ =0.0868, $p_{q2}$ =0.9270, $p_{q3}$ =0.0934, $p_{q4}$ =0.0000, $p_{q5}$ =0.0000, $p_{q6}$ =0.0021.
+We ask "are they the same?" ("yes" is the wrong answer) given the human-rated ground truth of being different patents.
+We gradually add scientific QA pairs from 1-6 for each patent to the original pair of patent texts, and then have Llama 3.1 8B re-judge.
+As shown in Figure 11, the baseline represents judge accuracy and confidence without any self-questioning or scientific retrieval.
+The placebo condition uses the same question but adds a randomly selected piece of text from the same paper as the found answers, matched in length to the actual answers.
+This is aiming for ablation studies on retrieval quality – what if the retrieved answer is of very low quality (even unrelated)?
+We observe that as more QAs from science are added, judge accuracy significantly increases and becomes better than placebo answers.
+The confidence of both correct and incorrect answers decreases, but the drop is significantly greater for incorrect answers.
+Interestingly, even the placebo answer — though merely a random, same-length excerpt from the paper — enhances understanding to some extent.
+However, unlike actual scientific answers, its effect does not scale with the number of questions, meaning there is some initial signal that helps judgment but LLM does not learn more over more rounds of relatively unrelated science.
+This echoes with the findings in the main paper: (1) the question alone already facilitates understanding, functioning as cues for better deployment of internal knowledge (we explored this further in Section 5.1), and (2) the placebo text still contains scientific content, which can contribute to comprehension.